### v1

# Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System

Abstract.
Many program transformations and optimizations require information about the future behavior of the program. A standard way to obtain this information is to build an intermediate program representation, then use a backwards program analysis to propagate relevant information against the flow of control back to the transformation/optimization site. We instead propose to use prophecy variables, which predict information about the future execution of the program, to enable such transformations and optimizations. While it is possible to use backwards analyses to obtain correct prophecy variable predictions, we instead propose to use repeated forward program execution to obtain correct prophecy variable predictions.
We implement prophecy variables in BuildIt, a lightweight domain specific language implementation system. BuildIt uses staged compilation to implement high performance domain specific languages embedded within a standard general purpose programming language (C++). The BuildIt first phase uses standard C++ program execution to generate optimized C, C++, and CUDA second phase code. This approach enables BuildIt to eliminate programming language implementation components such as parsers and intermediate representations, delivering a dramatic decrease in the engineering effort required to implement domain specific languages. The combination of prophecy variables and repeated forward program execution enables BuildIt to extend this approach to include transformations and optimizations that require information about the future execution of the program without backwards analyses and without the engineering overhead associated with implementing these analyses.
We formalize the use of prophecy variables for this purpose, discuss the implementation of prophecy variables and repeated execution in BuildIt, and present experimental results for BuildIt computations that benefit from optimizations enabled by the information that prophecy variables provide. To the best of our knowledge, we present the first implementation of prophecy variables for program analysis and optimization and the first use of repeated program execution to compute correct prophecy variable predictions.
copyright: none
1. Introduction
Over the last decade, domain specific languages (DSLs) have become a vehicle of choice for high-performance computations. Focusing on a specific domain enhances programmability by providing domain concepts directly in the programming language and enhances performance by enabling optimization techniques that leverage the focused domain to apply program transformations and optimizations. In comparison with alternative approaches that encapsulate high performance implementations of widely used operations such as matrix multiply in libraries, then structure the computation as a sequence of library calls, implementations of DSLs can implement optimizations that target the entire multioperation computation. Examples of such optimizations include efficiently managing data placement and movement across the full multioperation computation (as opposed to within a single operation), fusing sequences of operations into a single efficient integrated computation, and simplifying and specializing operators to specific values that occur at run time.
A common issue, however, is implementation overhead. DSL implementations often involve programming language implementation components such as parsers, intermediate representations, semantic checkers, and program analyses over a range of intermediate representations. The resulting implementation effort can require the language implementer to develop multiple tens of thousands of lines of code (in some cases over a hundred thousand lines of code, see Table 2), which can comprise a substantial obstacle to successful DSL development and deployment.
Standard DSL Implementation vs. Implementation Lines of Code BuildIt Speedup
BuildIt Standard DSL BuildIt Geometric Mean, Max, Min
GraphIt (g2) / GraphIt (BuildIt)(buildsl) 23,783 2,021 3%, 33%, -8%
COLA (shim) / SHIM (BuildIt) (shim) 7,500 5,700 1.4%, 32%, -30%
PCRE2(pcre2) / BREze (BuildIt) (tamara:meng-thesis:2023) 133,995 1,564 800%, 9100%, -19%
RE2(re2) / BREze (BuildIt) (tamara:meng-thesis:2023) 26,587 1,564 37%, 600%, -129%
Hyperscan(hyperscan) / BREze (BuildIt) (tamara:meng-thesis:2023) 187,033 1,564 -127%, 2000%, -1700%
Table 1. Implementation Effort and Performance of Standard DSL Implementations vs. Corresponding BuildIt Implementations. The first column (Standard DSL Implementation vs. BuildIt) presents the standard DSL implementation/BuildIt implementation of the same DSL. The next two columns (Implementation Lines of Code) present the number of lines of code in the standard DSL implementation vs. the number of lines of code in the corresponding BuildIt implementation. These numbers show that BuildIt implementations are typically much smaller (up to 100 times smaller) than standard implementations of the same DSL. The final column (BuildIt Speedup) presents the speedup of the BuildIt generated code relative to the code generated by the corresponding standard DSL implementation. Each DSL comes with a benchmark suite. We report three numbers for each DSL: 1) the geometric mean of the BuildIt speedup over all benchmarks, 2) the maximum BuildIt speedup over all benchmarks, and 3) the minimum BuildIt speedup over all benchmarks. These numbers show that BuildIt delivers performance that is at least comparable to and often significantly better than the performance provided by standard DSL implementations.
Component BuildIt(buildsl) GraphIt(g2)
Frontend - 9,593
Scheduling Language 151 2,401
Midend Analysis and Transformations - 9,601
Types and Operator Implementation 1,320 -
Code Generation 550 2,188
Total 2,021 23,783
Table 2. Implementation Lines of Code Breakdown for the GraphIt DSL. Each row presents the number of lines of code for a different implementation component for BuildIt and for the standard GraphIt DSL implementation. The first column (Component) identifies the component. The second column (BuildIt) presents the number of lines of code for the component in the BuildIt implementation. We note that the lightweight BuildIt approach eliminates the frontend and midend analysis and transformation components, leaving the BuildIt runtime as the largest component of the BuildIt implementation. The third column (GraphIt Compiler) presents the number of lines of code for the corresponding component of the standard GraphIt DSL implementation. These numbers highlight the implementation overhead associated with standard programming language implementation components such as frontend parsers and intermediate representation construction and processing. The overall result is that, for this DSL, BuildIt delivers an order of magnitude reduction in implementation effort in comparison with standard DSL implementation techniques.
1.1. BuildIt
In response to this situation, we developed BuildIt (buildsl; ajay-thesis), a system that leverages modern programming language features such as overloading to implement DSLs. Operating completely within a modern general purpose language, BuildIt implements compilation and optimization techniques that operate across the entire domain specific multioperation computation. BuildIt computations execute two stages. The first stage performs a combined partial evaluation, analysis, and compilation that generates optimized code . The second stage executes this optimized code (with the second stage execution typically comprising the overwhelming percentage of the overall computation). Because BuildIt is implemented as a collection of data structures implemented completely within a host general purpose language, it eliminates the need for language implementation components such as parsers and full blown intermediate representations. The result is a dramatic reduction in language implementation overhead. Table 2 quantifies this reduction by counting the number of lines of code in corresponding BuildIt and standard DSL implementations. Table 2 breaks down the lines of code for the GraphIt implementation components, highlighting the components that the lightweight BuildIt approach eliminates. These numbers show that BuildIt implementations can require up to 100 times fewer lines of code than standard implementations of the same DSL, a reduction substantial enough to transform the viability of many DSLs. And the performance results in Table 2 show that, in comparison with standard DSL implementations, BuildIt implementations typically exhibit comparable or better performance (buildsl; ajay-thesis).
1.2. Program Analysis in BuildIt
To apply the optimizations required to obtain acceptable performance, BuildIt must perform a range of program analyses. These take place during the BuildIt first phase and typically target the entire multioperation computation. For analyses that propagate information forward with the flow of control, BuildIt uses history variables (implemented as data structure fields in the underlying implementation language) to carry the information through the computation (OwickiG76; AbadiL91; rinard2020dataflow). These variables are assigned as the BuildIt first stage executes and carry the information through the first stage execution to be later accessed (again during the first stage) to optimize the generated second stage code.
1.3. Prophecy Variables, BuildIt, and Lightweight Program Analysis
Some optimizations, however, require information about the future program execution (history variables record information about the past, not future, execution). One standard way to obtain this information is to build an internal intermediate program representation, then use a backwards analysis to propagate information against the flow of control to reach the point where it is applied to optimize the generated code (DragonBook; appel2004modern; muchnick1997advanced; cooper2011engineering; kennedy2001optimizing). A drawback of attempting to apply this approach in the BuildIt context is the need to construct an appropriate program representation to enable the backwards analysis, an endeavor that would move the BuildIt implementation closer to standard DSLimplementations, increase the BuildIt implementation effort, and erode the implementation effort advantage that BuildIt currently exhibits over standard implementations of DSLs.
We instead present a novel use of prophecy variables to obtain information about the future program execution. Prophecy variables were initially developed to prove forward simulation relations in state machine refinement proofs that involve reasoning about future program behavior (AbadiL91). The basic idea behind prophecy variables in the BuildIt context is to use prophecy variable predictions to predict information about the future execution, then, when the information becomes available later in the execution, use prophecy variable preconditions to check that the prophecy variable correctly predicts the information.
In the refinement proof context, prophecy variable predictions are typically nondeterministic, with angelic nondeterminism enabling the refinement proof to go through if any of the nondeterministic predictions was correct (AbadiL91). In the BuildIt context, we instead use repeated program executions to discover correct prophecy variable predictions. When the BuildIt implementation encounters a violated prophecy variable precondition (this occurs when there is an incorrect prophecy variable prediction), it updates the prophecy variable to satisfy the precondition, propagates the update to the prophecy variable prediction point, terminates the BuildIt (first phase) execution, then reexecutes with the new prophecy variable prediction, repeating this process until all prophecy variable predictions are correct.
This mechanism enables BuildIt to use a standard general purpose language implementation, which provides only forward program execution, to implement transformations and optimizations that require information about future program executions. This mechanism preserves the implementation effort advantages that BuildIt exhibits over standard DSL implementations while supporting transformations and optimizations that normally require the implementation mechanisms and overheads traditionally associated with backwards analyses. And we note that prophecy variables are particularly well suited to the BuildIt staged approach — the repeated reexecutions that compute correct prophecy variable predictions all occur in the BuildIt first stage, with the reexecution overhead profitably amortized away by the generated second stage code, which typically performs the overwhelming majority of the overall computation.
1.4. Contributions
This paper makes the following contributions:
  • Program Analysis Information About Future Execution: It presents a novel technique, prophecy variable prediction via program reexecution, for obtaining program analysis information about the future execution of the program. This technique detects and corrects incorrect prophecy variable predictions, then reexecutes the program with the corrected prophecy variable predictions until all prophecy variable predictions are correct.
    A standard way to obtain program analysis information about the future execution is to build an intermediate program representation, then implement a backwards analysis that uses this intermediate representation to propagate information backwards against the flow of control.
    Prophecy variable prediction via reexecution, in contrast, reduces implementation effort because it requires no intermediate program representation, performs no backwards program analysis, and can be implemented directly on top of a standard programming language implementation, which provides only forward program execution.
  • Prophecy Variables in BuildIt: It presents the implementation of prophecy variables in the BuildIt system, which implements DSLs directly on top of a standard general purpose programming language without heavyweight language implementation components such as parsers or intermediate representations.
    Prophecy variable prediction via reexecution enables BuildIt to successfully obtain program analysis information about the future execution of the program without backwards analyses or the construction of an intermediate representation. The technique is therefore critical to enabling BuildIt to implement optimizations that require this information while maintaining the lightweight BuildIt DSL implementation approach.
    We note that the BuildIt prophecy and history variable mechanisms are encapsulated in the BuildIt data structures that implement the relevant BuildIt DSL, enabling application developers to write client code with little to no required knowledge of prophecy or history concepts.
  • Performance Results: It presents performance results that characterize the impact of two optimizations, CPU to GPU tensor preloading and convolution/ReLU fusion, that prophecy variables enable. The performance improvements due to these optimizations highlight the benefits that prophecy variables can bring in this context. It also presents results that characterize the implementation effort required to use prophecy variables in the benchmark DSLs.
The remainder of the paper is structured as follows. Section 2 presents an example that illustrates the use of prophecy variables in BuildIt. Section 3 presents a formal treatment of prophecy variables and correct prophecy variable prediction via program reexecution in the context of a core programming language. Section 4 discusses the implementation of prophecy variables in the BuildIt system and presents performance and engineering effort results for benchmark BuildIt computations that use optimizations enabled by prophecy variables. We discuss related work in Section 5 and conclude in Section 6.
2. Example
Figure 1 presents example BuildIt code for tensor computations using the Einstein summation notation (einstein_foundation_1916) (einsum notation) written in C++ (einsum-buildit). The example uses prophecy variables to predict tensors that may be accessed by a computation performed on a GPU. This information enables BuildIt to generate code that moves the accessed tensors to GPU memory before executing the GPU computation. The example uses several BuildIt types that guide the BuildIt staged execution. static_var and prophecy_var variables are first stage variables that exist only during the first stage (with the values of updated prophecy_var variables persisting across first stage reexecutions triggered by violated prophecy variable preconditions). dyn_var variables are second stage variables whose values are computed only during the second stage. Expressions and control flow that depend only on the values of static_var and prophecy_var variables are fully computed during the first stage. The first stage produces (optimized) second stage code that computes second stage expressions and control flow that depends on the values of dyn_var variables. Second stage expressions may contain first stage variables. BuildIt replaces these first stage variables in the generated second stage code with the corresponding first stage values that exist during first stage code generation, effectively specializing the generated second stage code to the first stage values.
The example defines a first stage variable current_execution_device (line 7 Figure 1) that records whether BuildIt is currently generating CPU or GPU code (the value of this variable is device_cpu when BuildIt generates code that executes on the CPU and device_gpu when BuildIt generates code that runs on the GPU).
The tensor data structure contains second stage fields m_buffer and m_gpu_buffer (lines 12-13 Figure 1) that store the tensor values on the CPU and GPU, respectively. The run_on_gpu(...) function (lines 37-58 Figure 1) takes a tensor computation f as a parameter and runs f on the GPU. Before the computation f executes, the generated second stage code copies the tensors that the computation f will read to the GPU memory (cudaMemcpyHostToDevice(...), line 42 Figure 1). A first stage loop iterates over all active tensors to identify the specific tensors that f may read and generate second stage code that copies these tensors to GPU memory (lines 38-44 Figure 1).
1enum device_type {
2 device_cpu,
3 device_gpu
4};
5
6// A global variable to track where we are currently running the code
7static_var<device_type> current_execution_device = device_cpu;
8
9template <typename T>
10struct tensor {
11 // Second stage variable to track the actual tensor values
12 dyn_var<T*> m_buffer;
13 dyn_var<T*> m_gpu_buffer;
14 static_var<std::vector<int>> m_sizes;
15
16 prophecy_var<TrueTop>* gpu_read = nullptr;
17 static_var<bool> gpu_written = false;
18
19 dyn_var<T> get_value(dyn_var<int> flat_index) {
20 if (current_execution_device == device_gpu) {
21 gpu_read->assert_requires(TrueTop::TRUE);
22 return m_gpu_buffer[flat_index];
23 }
24 return m_buffer[flat_index];
25 }
26 void operator=(const tensor& o) {
27 if (current_execution_device == device_gpu) {
28 gpu_written = true;
29 }
30 // Implementation to write to the Tensor
31 ...
32 }
33};
34
35// Function run tensor operations on GPU, block of code is
36// accepted as a lambda
37void run_on_gpu(std::function<void(void)> f) {
38 for (auto tensor: active_tensors) {
39 // Create a new prophecy variable with preferred value as False
40 tensor->gpu_read = new prophecy_var<TrueTop>(TrueTop::FALSE);
41 if (tensor->gpu_read->get_value() == TrueTop::TRUE) {
42 cudaMempcyHostToDevice(tensor->m_gpu_buffer, tensor->m_buffer, tensor->total_size());
43 }
44 }
45 current_execution_device = device_gpu;
46 // Actual implementation to move computation to GPU
47 buildit::dispatch_on_gpu(f);
48 current_execution_device = device_cpu;
49
50 for (auto tensor: active_tensors) {
51 if (tensor->gpu_written) {
52 cudaMemcpyDeviceToHost(tensor->m_buffer, tensor->m_gpu_buffer, tensor->total_size());
53 }
54 // Delete the prophecy variable
55 delete tensor->gpu_read;
56 tensor->gpu_written = false;
57 }
58}
59// Example program
60void matrix_mul(int M, int N, int O) {
61 tensor<float> A({M, N});
62 tensor<float> B({N, O});
63 tensor<float> C({M, O});
64 index i, j, k;
65 A[i][j] = 3.0;
66 B[i][j] = 4.0 + i + j;
67 run_on_gpu([&]() {
68 C[i][j] += A[i][k] * B[k][j];
69 });
70}
Figure 1. Prophecy Variables Predict Tensors Accessed by GPU Computations.
To enable run_on_gpu(...) to identify which tensors f will read, each tensor contains a boolean prophecy variable gpu_read (line 16 Figure 1) that is true if f will read the tensor. Together, the tensor prophecy variables implement a single conceptual prophecy variable that predicts the set of tensors that f will read, with the set implemented jointly by the boolean gpu_read prophecy variables in each tensor that indicate whether the tensor is a member of this set. These prophecy variables operate as follows:
  • Initialization: First stage code initializes the prophecy variables to false (line 40 Figure 1) — the default prediction is that the computation f will not read the tensor. The initial first stage execution therefore generates code that copies no tensors to the GPU memory.
  • Tensor Reads: Code in f invokes get_value(...) (lines 19-25 Figure 1) to read tensor values out of tensor memory. First stage code in get_value(...) uses the first stage current_execution_device variable (line 20 Figure 1) to determine whether the generated second stage code will execute on the GPU (and should therefore access values in m_gpu_buffer) or the CPU (and should therefore access m_buffer). In this way the generated second stage get_value(...) code is specialized for the specific device (GPU or CPU) on which it executes.
  • Prophecy Variable Analysis: To ensure that all tensors accessed by GPU code are correctly copied into GPU memory before the GPU code executes, the first stage code in get_value(...) (lines 20-21 Figure 1) implements a prophecy variable precondition
    (gpu_read->assert_requires(TrueTop::TRUE)) that requires the prophecy variable to correctly predict that GPU code may read the tensor.
    If the prediction is incorrect (and the value of the prophecy variable is false), BuildIt corrects the prediction (by setting the prophecy variable to true) and reexecutes the first stage computation with the new prophecy variable prediction. This first stage reexecution will now generate code that copies the corresponding tensor into GPU memory before the GPU computation executes (lines 41-43 Figure 1).
This mechanism generates potentially multiple first stage reexecutions triggered by prophecy variable mispredictions. The resulting prophecy variable corrections and reexecutions will eventually identify all (and only) those tensors that the generated GPU code may read, with the generated second stage CPU code correctly moving those tensors to GPU memory before executing the GPU computation.
We note that, unlike standard backwards analyses, this mechanism relies only on the ability to execute (and reexecute) the program written in a standard programming language. There is no intermediate representation construction and no analysis (backwards or forwards) of this (nonexistent) intermediate representation. The prophecy variable mechanism instead enables BuildIt to use standard forward program execution to correctly predict and move to GPU memory the tensors that arbitrary computations f may read.
We also note that tensor GPU computations often compute tensor results and these tensors should be transferred back to CPU memory when the GPU computation finishes. The code in Figure 1 uses the history variable gpu_written (line 17 Figure 1) for this purpose. First stage code in the operator = (lines 27-29 Figure 1) identifies which tensors the computation may write. First stage code in run_on_gpu(...) (lines 50-57 Figure 1) uses these history variables to identify and transfer (from GPU to CPU memory) tensors that the GPU computation may write. As this example illustrates, prophecy and history variables work together in the BuildIt system to enable the first stage computation to reason about both the future execution (prophecy variables) and past execution (history variables).
Finally, (as illustrated by the example program at lines 59-70 of Figure 1) we note that this entire prophecy variable mechanism is completely encapsulated in the implementation of the Tensor data structure. The application programmer does not need to understand prophecy variables, write prophecy variable code or even be aware of the existence of prophecy variables. This is consistent with the BuildIt approach, which uses mechanisms such as prophecy and history variables to encapsulate complex data management and optimization code inside data structure implementations instead of engineering a full blown DSL implementation using standard programming language implementation techniques.
3. Prophecy Variable Formalization
We next present and formalize the prophecy variables and prophecy variable analysis via reexecution.
3.1. Overview
We formalize prophecy variables and the prophecy variable analysis using a core programming language inspired by Glynn Winskell’s imperative language IMP (Winskel93). Notable differences include the introduction of labels for all commands and the use of variables $V$ instead of locations ($\mathbf{Loc}$(rinard2020dataflow). Specifically, the formalization works with programs $P$ that contain labeled commands of the form $l:c\in P$, where $l,g\in L$, $c\in C$. $\textbf{labels}(P)=\{l.l:c\in P\}$ is the set of labels in $P$. Labels are unique — no two labeled commands in $P$ have the same label $l$. An executing program operates on states $\sigma\in\Sigma$ that maps variables $v,w\in V$ to integer values $n,m\in N$. The standard operational semantics is modeled by a program execution transition relation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$. Execution starts at $\left\langle l_{0},\sigma_{0}\right\rangle$, where $l_{0}=\mathbf{first}(P)$ is the label of the first command to execute and $\sigma_{0}$ is the initial state. Execution terminates if it encounters an $l:\mathbf{halt}$ command.
We define the operational semantics by a set of program execution rules. Each rule starts with a program configuration $\left\langle l,\sigma\right\rangle$ to produce a next configuration $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$. Each rule has a set of preconditions that must be satisfied for the rule to execute. If the execution encounters an error, at least one of the relevant preconditions will not be satisfied and the execution will become stuck in a configuration $\left\langle l,\sigma\right\rangle$ such that $\left\langle l,\sigma\right\rangle\not\rightarrow$.
Each program analysis extends the state with a prophecy variable $\pi\in\Pi$, where $\left\langle\Pi,\leq\right\rangle$ is a lattice ordered by $\leq$ with least upper bound $\lor$ and greatest lower bound $\land$. The analysis also defines an extended operational semantics by extending the program execution rules to define an extended program execution relation $\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$. The extended program execution rules use $\pi$ to predict information about the future program execution, with incorrect predictions filtered out by new prophecy variable preconditions that check that the prediction was correct (with the execution becoming stuck if the prediction was not correct).
We require the extended program execution relation $\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$ not to introduce new executions. More precisely, we require the extended program execution relation to satisfy the following preservation property:
Definition 3.1.
(Preservation): If $\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$, then $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$.
Verifying the preservation property is typically straightforward as the updated program execution rules in the extended operational semantics typically have the same preconditions over $l$ and $\sigma$ and generate the same $l^{\prime}$ and $\sigma^{\prime}$ as the corresponding program execution rules in the standard operational semantics.
For each labeled command $l:c\in P$, there are two program points: $\bullet l$ (the program point before $l:c\in P$ executes) and $l\bullet$ (the program point after $l:c\in P$ executes). For each labeled command $l:c\in P$, the analysis produces an analysis result $\beta_{\bullet l}\in\Pi$ (drawn from same lattice $\left\langle\Pi,\leq\right\rangle$ as the prophecy variables $\pi$) at the program point $\bullet l$ before $l:c\in P$ executes.
Conceptually, the analysis is consistent with the extended operational semantics if it produces an analysis result that enables a corresponding transition in the extended operational semantics for each transition in the standard operational semantics. We formalize this requirement with the following Progress property:
Definition 3.2.
(Progress): If $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$, then $\left\langle l,\sigma,\beta_{\bullet l}\right\rangle\Rightarrow\left\langle l^{\prime},\sigma^{\prime},\beta_{\bullet l^{\prime}}\right\rangle$.
The progress property requires the analysis to produce correct prophecy variable predictions about all possible future executions (in the sense that the analysis results satisfy the prophecy variable preconditions that check for incorrect predictions).
We note that if Preservation (Definition 3.1) and Progress (Definition 3.2) both hold, then the relation $\sim$ defined by $\left\langle l,\sigma\right\rangle\sim\left\langle l,\sigma,\beta_{\bullet l}\right\rangle$ is a bisimulation relation (Milner89; Benthem96) between standard and extended program configurations (rinard2020dataflow).
3.2. Core Programming Language
$n,m\in N$ is the set of integers. $t\in T=\left\{\mathbf{true},\mathbf{false}\right\}$ is the set of truth values. Programs work with variables $v,w\in V$, arithmetic expressions $e\in E$, and boolean expressions $b\in B$ defined as follows:
$\begin{array}[]{rcl}E&::=&n|v|E_{0}+E_{1}|E_{0}-E_{1}|E_{0}\times E_{1}\\ B&::=&\mathbf{true}|\mathbf{false}|E_{0}=E_{1}|E_{0}\leq E_{1}|\mathbf{not}\;B|B_{0}\;\mathbf{and}\;B_{1}|B_{0}\;\mathbf{or}\;B_{1}\\ C&::=&\mathbf{skip}|v:=E|\mathbf{if}\;B\;\mathbf{then}\;g|\mathbf{goto}\;g|\mathbf{halt}|\mathbf{done}\end{array}$
$\mathbf{vars}(e)$ is the set of variables $v$ that $e$ reads, $\mathbf{vars}(b)$ is the set of variables $v$ that $b$ reads, and $\mathbf{vars}(c)$ is the set of variables that $c$ reads.
Each program $P$ is a sequence of labeled commands of the form $l:c$, where $l,g\in L$, $c\in C$. Given a program $P$ and label $l\in\mathbf{labels}(P)$, $l^{\prime}=\mathbf{next}(l)$ is the label $l^{\prime}$ of the next command (in the sequential execution order) in $P$ after $l:c\in P$. Conceptually, when the program executes a $l:\mathbf{halt}\in P$ command, the program stops executing in the $\mathbf{done}$ state. We therefore require if $l:\mathbf{halt}\in P$ and $l^{\prime}=\mathbf{next}(l)$, then $l^{\prime}:\mathbf{done}\in P$. We also require $\mathbf{next}(l)=g$ if $l:c=\mbox{if }l:\mathbf{goto}\;g\in P$ (but typically reference the branch target $g$ explicitly instead of $\mathbf{next}(l)$). We define the successors $\mathbf{succ}(l)$ of $l$ as follows:
[…]
and the predecessors $\mathbf{pred}(l)$ of $l$ as $\mathbf{pred}(l)=\left\{g.\mathbf{succ}(g)=l\right\}$.
3.3. Operational Semantics
A program in execution maintains a state $\sigma:V\rightarrow N\in\Sigma$, where $\sigma(v)$ is the value of the variable $v$ in state $\sigma$. Given an arithmetic expression $e$ and state $\sigma$, we define the arithmetic expression evaluation relation $\left\langle e,\sigma\right\rangle\rightarrow n$ as the smallest relation (under subset inclusion) over $E\times\Sigma\times N$ that satisfies the arithmetic expression evaluation rules in Figure 4. Given a boolean expression $b$ and state $\sigma$, we define the boolean expression evaluation relation $\left\langle b,\sigma\right\rangle\rightarrow t$ as the smallest relation (under subset inclusion) over $B\times\Sigma\times T$ that satisfies the boolean expression evaluation rules in Figure 4.
$\left\langle n,\sigma\right\rangle\rightarrow n$ $\left\langle v,\sigma\right\rangle\rightarrow\sigma(v)v\in\mathbf{dom}\;\sigma$ $\left\langle e_{0}+e_{1},\sigma\right\rangle\rightarrow n_{0}+n_{1}\left\langle e_{0},\sigma\right\rangle\rightarrow n_{0}\;\left\langle e_{1},\sigma\right\rangle\rightarrow n_{1}$
$\left\langle e_{0}-e_{1},\sigma\right\rangle\rightarrow n_{0}-n_{1}\left\langle e_{0},\sigma\right\rangle\rightarrow n_{0}\;\left\langle e_{1},\sigma\right\rangle\rightarrow n_{1}$ $\left\langle e_{0}*e_{1},\sigma\right\rangle\rightarrow n_{0}*n_{1}\left\langle e_{0},\sigma\right\rangle\rightarrow n_{0}\;\left\langle e_{1},\sigma\right\rangle\rightarrow n_{1}$
Figure 2. Arithmetic Expression Evaluation Rules
$\left\langle\mathbf{true},\sigma\right\rangle\rightarrow\mathbf{true}$ $\left\langle\mathbf{false},\sigma\right\rangle\rightarrow\mathbf{false}$ $\left\langle e_{0}=e_{1},\sigma\right\rangle\rightarrow n_{0}=n_{1}\left\langle e_{0},\sigma\right\rangle\rightarrow n_{0}\;\left\langle e_{1},\sigma\right\rangle\rightarrow n_{1}$
$\left\langle e_{0}\leq e_{1},\sigma\right\rangle\rightarrow n_{0}\leq n_{1}\left\langle e_{0},\sigma\right\rangle\rightarrow n_{0}\;\left\langle e_{1},\sigma\right\rangle\rightarrow n_{1}$ $\left\langle\mathbf{not}\;b,\sigma\right\rangle\rightarrow\mathbf{false}\left\langle b,\sigma\right\rangle\rightarrow\mathbf{true}$ $\left\langle\mathbf{not}\;b,\sigma\right\rangle\rightarrow\mathbf{true}\left\langle b,\sigma\right\rangle\rightarrow\mathbf{false}$
$\left\langle b_{0}\;\mathbf{and}\;b_{1},\sigma\right\rangle\rightarrow t_{0}\;\mathbf{and}\;t_{1}\left\langle b_{0},\sigma\right\rangle\rightarrow t_{0}\;\left\langle b_{1},\sigma\right\rangle\rightarrow t_{1}$ $\left\langle b_{0}\;\mathbf{or}\;b_{1},\sigma\right\rangle\rightarrow t_{0}\;\mathbf{or}\;t_{1}\left\langle b_{0},\sigma\right\rangle\rightarrow t_{0}\;\left\langle b_{1},\sigma\right\rangle\rightarrow t_{1}$
Figure 3. Boolean Expression Evaluation Rules
$\left\langle l,\sigma\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma[v\mapsto n]\right\rangle l:v:=e\;\in\;\mathbf{P}\;\;\left\langle a,\sigma\right\rangle\rightarrow n$
$\left\langle l,\sigma\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma\right\rangle l\;:\;\mathbf{if}\;b\;\mathbf{then}\;g\;\in\;\mathbf{P}\;\;\;\left\langle b,\sigma\right\rangle\rightarrow\mathbf{false}$ $\left\langle l,\sigma\right\rangle\rightarrow\left\langle g,\sigma\right\rangle l:\mathbf{if}\;b\;\mathbf{then}\;g\;\in\;\mathbf{P}\;\;\;\left\langle b,\sigma\right\rangle\rightarrow\mathbf{true}$
$\left\langle l,\sigma\right\rangle\rightarrow\left\langle g,\sigma\right\rangle l:\mathbf{goto}\;g\;\in\;\mathbf{P}$ $\left\langle l,\sigma\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma\right\rangle l:\mathbf{skip}\;\in\;\mathbf{P}$ $\left\langle l,\sigma\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma\right\rangle l:\mathbf{halt}\in P\;\;\;\mathbf{next}(l):\mathbf{done}\in P$
Figure 4. Program Execution Rules
Given a program $P$, the operational semantics works with configurations of the form $\left\langle l,\sigma\right\rangle$, where $l$ is the label of a labeled command $l:c\in P$ and $\sigma:V\rightarrow N$ is an environment that maps variables $v\in V$ to values $n\in N$. We define the program execution relation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$ as the smallest relation (under subset inclusion) over $L\times\Sigma\times L\times\Sigma$ that satisfies the program execution rules in Figure 4.
Given a program $P$, a sequence of configurations $s=\left\langle l_{0},\sigma_{0}\right\rangle\rightarrow\cdots\rightarrow\left\langle l_{n},\sigma_{n}\right\rangle\cdots$, where $l_{0}=\mathbf{first}(P)$ and $\sigma_{0}=\emptyset$ is the initial state of $P$, is an execution of $P$. If $s$ is a finite sequence of length $n+1$, then $s$ is a finite execution of $P$ and $\left\langle l_{n},\sigma_{n}\right\rangle$ is the last element in the sequence. If $l_{n}:\mathbf{done}\in P$, then the sequence is a complete execution of $P$, otherwise, it is a partial execution of $P$. If $s$ is a partial execution of $P$ and $\left\langle l_{n},\sigma_{n}\right\rangle\not\rightarrow$, then $s$ is a stuck execution of $P$ (an execution of $P$ can become stuck if the precondition of an expression evaluation or command execution rule is not satisfied). If $s$ is an infinite sequence, then the sequence is an infinite execution of $P$. The $\mathbf{traces}(P,\rightarrow)$ are all of the complete and infinite executions of $P$.
3.4. Extended Operational Semantics With Prophecy Variables
We extend the expression evaluation and program execution relations with prophecy variables $\pi$ to obtain an extended operational semantics defined by relations $\left\langle e,\sigma,\pi\right\rangle\rightarrow n$, $\left\langle b,\sigma,\pi\right\rangle\rightarrow t$, and $\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$. We focus on analyses that work with sets of items, so that the prophecy variables $\pi$ are sets of items. We extend the preconditions of the program execution rules (Figure 4) with a (nondeterministic) prophecy variable prediction of the form $\pi^{\prime}\subseteq\pi\cup S$, where $S$ is a set of items (note that often $S=\emptyset$, in which case the prediction is $\pi^{\prime}\subseteq\pi$). We also extend the preconditions of expression evaluation and program execution rules (Figures 4, 4, and 4) with an (optional) prophecy variable precondition of the form $S\subseteq\pi$, which requires the prophecy variable $\pi$ to contain a set of predicted items $S$. If an incorrect prediction causes the prophecy variable to not contain $S$, the execution becomes stuck. The verification of the Preservation property is immediate — the expression evaluation and program execution rules in the extended operational semantics have the same preconditions over $l$ and $\sigma$ and generate the same $l^{\prime}$ and $\sigma^{\prime}$ as the corresponding program execution rules in the standard operational semantics.
$\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle\mathbf{next}(l),\sigma[v\mapsto n],\pi^{\prime}\right\rangle l:v:=e\;\in\;\mathbf{P}\;\;\left\langle a,\sigma,\pi\right\rangle\Rightarrow n\;\;\;\pi^{\prime}\subseteq\pi\cup\left\{v\right\}$
$\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle\mathbf{next}(l),\sigma,\pi^{\prime}\right\rangle l:\mathbf{if}\;b\;\mathbf{then}\;g\;\in\;\mathbf{P}\;\;\;\left\langle b,\sigma,\pi\right\rangle\Rightarrow\mathbf{false}\;\;\;\pi^{\prime}\subseteq\pi$ $\left\langle l,\sigma,\pi\right\rangle\Rightarrow\left\langle g,\sigma,\pi^{\prime}\right\rangle l:\mathbf{if}\;b\;\mathbf{then}\;g\;\in\;\mathbf{P}\;\;\;\left\langle b,\sigma,\pi\right\rangle\Rightarrow\mathbf{true}\;\;\;\pi^{\prime}\subseteq\pi$
$\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle g,\sigma,\pi^{\prime}\right\rangle l:\mathbf{goto}\;g\;\in\;\mathbf{P}\;\;\;\pi^{\prime}\subseteq\pi$ $\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma,\pi^{\prime}\right\rangle l:\mathbf{skip}\;\in\;\mathbf{P}\;\;\;\pi^{\prime}\subseteq\pi$ $\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle\mathbf{next}(l),\sigma,\pi^{\prime}\right\rangle l:\mathbf{halt}\in P\;\;\;\mathbf{next}(l):\mathbf{done}\in P\;\;\;\pi^{\prime}\subseteq\pi$
Figure 5. Program Execution Rules for Live Variable Prediction with Prophecy Variables
For example, the program execution rules in Figure 5 instantiate this framework to predict live variables (variables that may be read before they are written in the future program execution): The rule for $l:v:=e\in P$ uses the prophecy variable prediction $\pi^{\prime}\subseteq\pi\cup\left\{v\right\}$ to predict that 1) $v$ may become live after $l:v:=e\in P$ and 2) some subset of $\pi\cup\left\{v\right\}$ (including $v$) may become dead after $l:v:=e\in P$. Variables may become dead either because $e$ contained the last read to the variable before the variable is reassigned or because $v$ itself is not read before it is reassigned. Similarly, the rule for $l:\mathbf{if}\;b\;\mathbf{then}\;g\in P$ uses the prophecy variable prediction $\pi^{\prime}\subseteq\pi$ to predict that some subset of the predicted live variables $\pi$ before $l:\mathbf{if}\;b\;\mathbf{then}\;g\in P$ will no longer be live after $l$, for example because $b$ contained the last access to a variable before the variable is reassigned. The remaining program execution rules use the prophecy variable prediction $\pi^{\prime}\subseteq\pi$ — because their commands all have a single successor, the prophecy variable analysis (Section 3.5) will compute a least (under subset inclusion) analysis result in which $\pi^{\prime}=\pi$.
Of course, it is possible for the program execution rules to mispredict which variables will become dead after executing $l:v:=e\in P$ or $l:\mathbf{if}\;b\;\mathbf{then}\;g\in P$. The extended operational semantics therefore updates the variable read rule to include the prophecy variable precondition $\left\{v\right\}\subseteq\pi$, which requires that every variable $v$ read during expression evaluation must be predicted live. With this precondition, all executions that mispredict a live variable become stuck at the command that attempts to read the mispredicted variable. Here $\mathbf{dom}\;\sigma$ is the domain of $\sigma$ viewed as a function — the set of variables $v$ for which $\sigma(v)$ is defined.
$\left\langle v,\sigma,\pi\right\rangle\Rightarrow\sigma(v)v\in\mathbf{dom}\;\sigma\;\;\;\left\{v\right\}\subseteq\pi$
3.5. Prophecy Variable Analysis via Repeated Program Execution
It is possible to use a backwards dataflow analyses to obtain prophecy variable values that correctly predict future program executions (rinard2020dataflow). These analyses, however, work with an explicit program representation to propagate information backwards against the flow of control. They are therefore incompatible with the BuildIt staged compilation approach, which executes within a standard forward program execution environment to eliminate language implementation components such as intermediate representations. We therefore use an alternate approach that uses repeated program reexecutions to compute prophecy variable values.
Algorithms 1, 2, and 3 implement this program reexecution based prophecy variable analysis. Together they produce prophecy variable analysis results $\beta_{\bullet l}$ at each program point $\bullet l$. After initializing all $\beta_{\bullet l}$ to $\emptyset$, Algorithm 1 executes the program using the standard semantics to check the prophecy variable preconditions (lines 6-12 Algorithm 1). Here we use the notation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$ with $S_{1}\subseteq\pi$ (line 6 Algorithm 1) to indicate that the extended program transformation rule that generates the corresponding $\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$ transition has the prophecy variable precondition $S_{1}\subseteq\pi$. If the algorithm encounters a violated prophecy variable precondition (line 7 Algorithm 1), it updates the prophecy variable ($\beta_{\bullet l}$) to satisfy the precondition (line 8 Algorithm 1), invokes Algorithm 2 to ensure that the relevant prophecy variable prediction constraints hold, then throws a ProphecyVariableMisprediction exception to trigger a reexecution with the new prophecy variable values.
Algorithm 1 Execute
1:
2:for all $l\in\mathbf{labels}(P)$ do $\beta_{\bullet l}\leftarrow\emptyset$
3:end for
[…]
5:procedure Execute($\left\langle l,\sigma\right\rangle$)
6:  for all $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$ with $S_{1}\subseteq\pi$ do
7:   if $S_{1}\not\subseteq\beta_{\bullet l}$ then
8:     $\beta_{\bullet l}\leftarrow\beta_{\bullet l}\cup S_{1}$
9:     Solve($l$)
10:     throw ProphecyVariableMisprediction
11:   end if
12:  end for
13:  for all $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$ with $\pi^{\prime}\subseteq\pi\cup S_{2}$ do
14:   $C\leftarrow C\cup\left\{\beta_{\bullet l^{\prime}}\subseteq\beta_{\bullet l}\cup S_{2}\right\}$
15:   Execute($\left\langle l^{\prime},\sigma^{\prime}\right\rangle$)
16:  end for
17:end procedure
Algorithm 2 Solve
1:procedure Solve($l$)
2:  for all $\beta_{\bullet l}\subseteq\beta_{\bullet l^{\prime}}\cup S_{2}\in C$ such that $\beta_{\bullet l}\not\subseteq\beta_{\bullet l^{\prime}}\cup S_{2}$ do
3:   $\beta_{\bullet l^{\prime}}\leftarrow\beta_{\bullet l^{\prime}}\cup(\beta_{\bullet l}-S_{2})$
4:   Solve($l^{\prime}$)
5:  end for
6:end procedure
Algorithm 3 Driver
1:procedure Driver
2:  try Execute
3:  catch (ProphecyVariableMisprediction) Execute
4:end procedure
The loop in lines 13-16 Algorithm 1 collects encountered prophecy variable predictions into a set of prophecy variable prediction constraints $C$. These constraints will be used by Algorithm 2 to update the prophecy variables to ensure that the prophecy variable predictions hold after a prophecy variable update. Here we use the notation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$ with $\pi^{\prime}\subseteq\pi\cup S_{2}$ to indicate that the extended program transformation rule that generates the corresponding $\left\langle l,\sigma,\pi\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime},\pi^{\prime}\right\rangle$ transition has the prophecy variable prediction $\pi^{\prime}\subseteq\pi\cup S_{2}$.
Finally, Algorithm 3 repeatedly reexecutes the program whenever Algorithm 1 encounters a prophecy variable misprediction, terminating when all prophecy variable preconditions hold.
The algorithm produces the least (under subset inclusion) prophecy variable analysis results $\beta_{\bullet l}$ that satisfy both the prophecy variable preconditions and the prophecy variable prediction constraints. These $\beta_{\bullet l}$ values therefore ensure that no program execution will become stuck because of an incorrect prophecy variable prediction, ensuring the Progress property.
4. Applications
We next present the use of BuildIt prophecy and history variables for optimizing computations written in a BuildIt DSL for performing einsum tensor computation and a BuildIt DSL for constructing neural networks. We present an overview of BuildIt (Section 4.1), prophecy variables in BuildIt (Section 4.2) and then the two DSLs along with a benchmark application in each DSL (Sections 4.3 and 4.4). We include performance results and development effort statistics for each of the DSL application combinations.
4.1. BuildIt Overview
BuildIt is a lightweight DSL implementation system (ajay-thesis). BuildIt programs execute two stages. The first stage executes as standard C++ code that generates optimized C, C++, and/or CUDA code that executes in the second stage. In a typical BuildIt program the vast majority of the computation is performed by the second stage.
The two stage execution is controlled by two type templates static_var<T> and dyn_var<T>. As illustrated by the example in Section 2, the first stage fully evaluates all expressions and control flow of type static_var<T>, with the generated second stage code evaluating all expressions of type dyn_var<T>. BuildIt supports arbitrarily complex control flow over both static_var<T> and dyn_var<T> expressions.
BuildIt is implemented as C++ data structures that rely only on the standard C++ execution environment. These data structures encapsulate BuildIt concepts, constructs, optimizations, transformations, and code generation steps, enabling application developers to write straightforward client code with little to no required knowledge of the details of BuildIt prophecy or history variable concepts or constructs.
BuildIt generates optimized second stage code specialized for the specific first phase values. The BuildIt first stage code generation often uses history variables to propagate information about the past execution of the program to first stage optimization or code generation sites. An example of such a history variable is the gpu_written variable in the example code from Section 2 (lines 17 and 51-53 in Figure 1). As this example shows, the BuildIt static_var<T> type directly supports the implementation of such history variables.
4.2. Prophecy Variables in BuildIt
We extend BuildIt to introduce a new type-template, prophecy<T>. Like static_var<T> variables, prophecy<T> variables are completely evaluated in the first stage with the exception that the values of updated prophecy<T> variables persist across reexecutions. prophecy<T> also has the constraint that the wrapped type T must be a lattice and that all updates to prophecy variables implement lattice joins. static_var<T> variables can be used to record and propagate first stage information about the past progam execution to program transformation and optimization sites; prophecy<T> variables can be used at program transformation and optimization sites to predict information about the future program execution. Both history and prophecy variables can be used to conditionally generate different code. Together, static_var<T> and prophecy<T> variables can be used to collect information about both the previous and future execution of the program, then transfer that information to relevant transformation and optimization sites, eliminating the need for traditional program analysis algorithms that operate on intermediate program representations.
BuildIt also uses program reexecution to explore and generate control flow that involves dyn_var<T> expressions (buildsl; ajay-thesis). We repurpose this program reexecution capability to implement the prophecy variable analysis formalized in Section 3.
4.3. Tensor Data Movement for GPU Kernels
This benchmark involves a BuildIt DSL that can evaluate arbitrary multidimensional tensor operations specified as einsum expressions (einsum-buildit). In this DSL the the dimensions and sizes of each tensor are first stage values. As a BuildIt DSL, the DSL is implemented as a C++ data structure library. The first stage executes with the standard C++ execution environment to generate second stage tensor code, including the generation of CUDA code for GPU kernels. The library introduces the key type Tensor<E>, where the first stage template parameter E is the element type of the tensor. The first stage tensor dimensions are stored internally as static_var<std::vector<int>> members. The second stage tensor values are stored in members of type dyn_var<E*>. Tensor operations such as [], +, *, =, += and *= are overloaded to implement the einsum operations.
Figure 6 presents a matrix multiply program written in the BuildIt einsum DSL, with the DSL run_on_gpu(...) operation marking the matrix multiply operation C[i][j] = A[i][k]*B[k][j] for generation as a GPU kernel. As this program illustrates, the BuildIt tensor library encapsulates all of the BuildIt mechanism required to generate optimized code, enabling developers to operate effectively without knowledge or even awareness of prophecy variables or other BuildIt mechanisms.
Before the generated second stage code launches the kernel, buffers to hold the accessed tensors must be allocated on the GPU and moved from the CPU to the GPU. A complication is that generating the tensor movement code requires information about the future execution of the program (specfically, which tensors the kernel will access when it runs). As outlined in Section 2, prophecy variables encapsulated in the Tensor<E> data structure enable the BuildIt first stage to generate second stage code that moves the read tensors to the GPU when the first stage enters run_on_gpu(...), conceptually before the first stage processes the kernel code, and without the construction of an intermediate representation or static analysis of any (nonexistent) intermediate representation.
The example tensor code in Figure 1 allocates buffers for all tensors on both the CPU and GPU. Some tensors, however, may never be resident on the GPU at all. The BuildIt einsum DSL handles this situation by adding two new prophecy variables to the BuildIt Tensor<E> data structure — a boolean needs_gpu (which is true if the tensor is ever accessed by the GPU) and a boolean gpu_read (which is true if a specific GPU kernel reads the tensor).
Prophecy variable preconditions in tensor operations that execute on the GPU require the prophecy variables to correctly predict 1) tensors that may participate in GPU computations (the needs_gpu prophecy variable) and 2) the tensors that a given operation will read (the gpu_read prophecy variable). The Tensor<E> constructor uses the needs_gpu prophecy variable to determine whether to allocate a GPU buffer for the tensor; run_on_gpu(...) uses the gpu_read prophecy variables to determine which tensors to move to the GPU before executing the kernel as in Section 2. We consider the prophecy variable implementation overhead to be very reasonable — the DSL is implemented in 608 lines of C++ code; 25 of these 608 lines of code deal with prophecy variables.
1void matrix_mul(int M, int N, int O) {
2 tensor<float> A({M, N});
3 tensor<float> B({N, O});
4 tensor<float> C({M, O});
5 index i, j, k;
6 A[i][j] = 3.0;
7 B[i][j] = 4.0 + i + j;
8 run_on_gpu([&]() {
9 C[i][j] += A[i][k] * B[k][j];
10 });
11}
Figure 6. Matrix-Matrix Multiplication in the Einsum DSL (also see lines 60-70 Figure 1). The two inputs and the outputs are of sizes M x N, N x O and M x O respectively where M, N and O are first stage parameters.
We present performance results for two computations: 1) matrix-matrix multiply (six tensors, three of which are ever resident on the GPU) and 2) matrix-vector multiply (two matrices and four vectors, with one matrix and one vector ever resident on the GPU). We consider two alternative GPU tensor movement implementations: 1) moving all tensors to the GPU before executing the GPU kernel (we acknowledge that this strategy becomes infeasible for applications that manipulate large numbers of tensors) and 2) using a runtime data movement strategy ( CUDA unified memory, which uses a page fault based mechanism to move data between CPU and GPU (Harris2017UnifiedMemory)). Table 8 presents the minimum execution time in microseconds for GPU matrix-matrix multiplication running with the three strategies and four matrix sizes across ten runs. Table 8 presents the minimum execution time in microseconds for GPU matrix-vector multiplication running with the three strategies and four matrix sizes across ten runs. The hardware platform for these experiments is an NVIDIA DGX-1 with an Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz and a Tesla V100-SXM2-32GB GPU. BuildIt with prophecy variables always exhibits the best performance, with the relative performance gap narrowing as the matrix size increases. We attribute this phenomenon to the fact that, because matrix multiply performs $O(n^{3})$ computation and $O(n^{2})$ communication, at larger matrix sizes the computation time increasingly dominates so that the communication mechanism plays a smaller role in the overall performance.
Size Copy Unified Prophecy
All Memory Enabled
64x64 149 239 43
128x128 284 269 94
256x256 840 615 248
512x512 3053 1647 1059
1024x1024 9,863 6,343 4,244
2048x2048 46,552 32,395 26,803
4096x4096 380378 316443 307,830
Figure 7. Execution time in microseconds for Matrix-Matrix Multiplication with the three strategies - copying all tensors to GPU memory, using unified memory, and BuildIt with prophecy variables, which copies only accessed tensors to GPU memory. We report minimum execution times in microseconds across 10 runs.
    Size Copy Unified Prophecy All Memory Enabled 64x64 112 79 35 128x128 173 132 60 256x256 362 237 129 512x512 1,199 621 468 1024x1024 3,061 1,356 1,054 2048x2048 9,647 3,564 2,854 4096x4096 34,341 9,754 9,199 Figure 8. Execution time in microseconds for Matrix-Vector Mulitiplication with the three strategies - copying all tensors to GPU memory, using unified memory, and BuildIt with prophecy variables, which copies only accessed tensors to GPU memory. We report minimum executions times in microseconds across 10 runs.
Table 9 presents the number of first stage program executions required to 1) fully explore all possible dynamic control flow paths (this is a baseline BuildIt technique for ensuring correct second stage code generation (buildit; ajay-thesis)) and 2) obtain correct prophecy variable values. It also presents the first stage execution times. Note that the use of prophecy variables can increase the number of dynamic control flow paths that BuildIt must explore to generate correct second stage code, which in turn causes an increase in the number of dynamic control flow exploration executions with prophecy variables and a corresponding increase in first stage execution times.
Matrix-Matrix Multiply Copy All Unified Memory Prophecy Variable
Prophecy Variable Related Runs - - 7
Dynamic Control Flow Runs 25 25 113
First Stage Execution Time 9,212 us 8,336 us 38,335 us
Matrix-Vector Multiply Copy All Unified Memory Prophecy Variable
Prophecy Variable Related Runs - - 7
Dynamic Control Flow Runs 19 19 89
First Stage Execution Time 7,154 us 6,180 us 27,167 us
Figure 9. Number of first stage program executions and first stage execution times for Matrix Matrix Multiply and Matrix Vector Multiply. We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
4.4. Neural Network Operation Fusion
This benchmark involves a machine learning DSL that implements neural network operations such as convolutions, fully connected layers, and activation functions including ReLU. The DSL allows developers to specify complete networks by specifying a sequence of these operations, with BuildIt generating second stage code that implements the resulting model.
Figure 10 presents an example BuildIt program in this DSL. The program performs a convolution operation followed by a ReLU operation (note that the ReLU operation takes an activation threshold as a parameter). An unoptimized implementation traverses the tensor once for the convolution, then again for the ReLU. But because the ReLU is a pointwise operation, it can be directly fused into the implementation of the convolution, eliminating the second iteration over the tensor to compute the ReLU. We note that this kind of operation fusion is an important optimization for obtaining good performance on neural network computations in general.
In the BuildIt implementation of this DSL, history and prophecy variables work together to enable BuildIt to fuse a ReLU into the generated second stage convolution code. Control flow complicates the correctness considerations associated with applying this operation fusion: the next operation after the convolution may be a ReLU on some control flow paths but not others (in which case BuildIt should not be fuse a ReLU into the convolution). Even if the next operation is a ReLU on all control flow paths, all of these ReLUs must have the same activation threshold for BuildIt to fuse the ReLU into the convolution. The BuildIt implementation deals with these complications with the following history and prophecy variables:
  • History Variable: A history variable is_last_convolution tracks the whether last performed tensor operation is a convolution.
  • Prophecy Variables: A prophecy variable is_next_relu predicts whether 1) all operations that immediately follow a convolution are ReLU operations and 2) if so, whether all of these ReLU operations have the same activation threshold (in which case the prophecy variable also stores the activation threshold).
The tensor operations work with these history and prophecy variables as follows. The corresponding code is all first stage code:
  • Operation Tracking: At the end of each invoked tensor operation, the implementation updates the history variable to identify that this tensor operation was the last performed tensor operation.
  • Immediately Following Operations: Each invoked tensor operation checks the history variable upon entry to determine if the last performed operation was a convolution. If so, it requires the prophecy variable to correctly reflect the fact that the invoked operation may immediately follow the convolution. If the invoked operation is a ReLU, the prophecy variable must also correctly reflect the ReLU activation threshold for that ReLU.
  • Optimization: The convolution operation checks the prophecy variable to determine if all operations immediately following the convolution are ReLU operations with the same activation threshold. If so, it generates second stage code that fuses the ReLU into the convolution. Otherwise it generates the convolution code only.
    The ReLU operation checks the prophecy variable to determine if all immediately following operations are ReLU operations with the same activation threshold. If so, it skips the ReLU (the first phase ReLU generates no second stage code) because the ReLU was fused into the convolution operation. Otherwise it generates code to implement the ReLU operation.
We consider the prophecy variable implementation overhead to be reasonable — the DSL is implemented in 161 lines of C++ code; 48 of these 161 lines of code deal with prophecy or history variables.
1void network(void) {
2 ml::tensor input(102400, {...});
3 ml::tensor weight(9, {...});
4
5 ml::tensor conv_out = ml::convolve(input, weight);
6 ml::tensor out = ml::relu(conv_out, 1.35);
7 ... // downstream operations
8}
Figure 10. BuildIt Code for Convolution Followed by ReLU.
Image Fusion Fusion
Size Disabled Enabled
10240 $\times$ 3 26 25
102400 $\times$ 3 267 243
1024000 $\times$ 3 2,594 2,341
10240 $\times$ 9 76 73
102400 $\times$ 9 770 742
1024000 $\times$ 9 7,953 7,451
Figure 11. Execution times in microseconds for 1-D convolution followed by ReLU with fusion disabled and enabled for different input and filter sizes.
    Convolution-ReLU Fusion Fusion Disabled Enabled Prophecy Variable Runs - 4 Dynamic Control Flow Runs 33 72 First Stage Execution Time 9,872 us 14,701 us Figure 12. Number of first stage program executions and first stage execution times. We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
We report execution times for the benchmark program in Figure 10 with the convolution/ReLU fusion optimization disabled and enabled. Table 12 presents execution times in microseconds for a 1-D convolution followed by a ReLU for a range of input sizes (10240 to 1024000) and filter sizes (9 point and 3 point filters). The hardware platform for these experiments is a 12th Gen Intel(R) Core(TM) i9-12900HK with 32GB of memory. Table 12 presents the number of first stage program executions and execution times required to fully explore all possible dynamic control flow paths and obtain correct prophecy variable values.
We note that neural networks typically consist of multiple layers including activation functions and other operations (such as normalization). Together, history and prophecy variables enable BuildIt to fully support optimizations that operate across layers and operations to generate efficient second stage code that fuses operations to eliminate unnecessary intermediate temporaries and repeated tensor traversals.
5. Related Work
We survey related work in prophecy and history variables and implementation techniques for domain specific languages. To the best of our knowledge, our research is the first to demonstrate the use of prophecy variables for program analysis in any programming language implementation and the first to propose program reexecution to obtain correct prophecy variable values.
5.1. Prophecy and History Variables
Many of the concepts that appear in simulation relation proofs for state machines also appear in the program verification, dataflow analysis, and abstract interpretation literature. For example, history variables were first introduced in the program verification literature (OwickiG76), abstraction functions, originally introduced in the program verification literature (Hoare72), can be seen as a form of refinement mappings, and program analyses can be seen as establishing a simulation relation between an abstract interpretation of the program (which plays the role of the specification) and concrete executions of the program (which play the role of the implementation) (CousotC77; CousotC92). It is also known that, in this context, backward or reverse simulation relations can be used to establish the correspondence between backward analyses (which extract information about the future execution) and program executions (CousotC92; SchmidtS98).
A recent (unpublished in any peer reviewed venue) manuscript introduces prophecy variables to enable forward reasoning about program analysis properties that involve the future execution of the program (rinard2020dataflow). To the best of our knowledge, this manuscript is the first to introduce prophecy variables for program analysis (here we contrast with the use of prophecy variables for program verification (JungLPRTDJ20; Vafeiadis08; ZhangFFSL12) as well as the traditional use of prophecy variables for proving forward simulation relations between state machines (AbadiL91)). The manuscript uses prophecy variables to specify two program analysis problems, live variables and very busy expressions, that require information about the future execution of the program and uses a backwards dataflow analysis to obtain correct prophecy variable predictions.
While we build on the formalism presented in this manuscript, our research differs in that 1) we use repeated forward program executions, not backwards dataflow analysis, to find correct prophecy variable predictions, 2) we implement prophecy variables in the BuildIt staged DSL implementation system ( (rinard2020dataflow) presents no implementation), 3) we show how prophecy variables enable the BuildIt first stage, which uses only standard forward program execution, to obtain and exploit information about the future execution of the program, and 4) we report results from BuildIt prophecy variable implementations. The overall result is the extension of the scope of the lightweight BuildIt implementation approach to include optimizations (such as those presented in Section 4) that require information about the future execution of the program.
Cobalt enables compiler developers to specify a range of dataflow optimizations (such as constant propagation and partial dead assignment elimination) (LernerMC03). Each optimization is specified by a transformation pattern whose guard specifies a condition over sequences of actions in paths in the program representation that must hold for the transformation to be legal. Cobalt has separate constructs for specifying forward and backward optimizations — forward guards reason about forward properties, backward guards reason about backward properties.
Rhodium was developed to eliminate the complexity of using temporal logic formulas and implicit dataflow facts to reason about paths in control-flow graphs (LernerMRC05). Rhodium uses de facto abstraction functions (expressed as predicates over concrete program states) to establish the connection between concrete program states and dataflow facts and state extensions (a form of instrumented semantics) to support analyses that extract information about the past execution of the program not present in standard concrete program states. Like Cobalt, Rhodium has separate support for forward and backward analyses; subsequent work on automatically inferring correct propagation rules supports only forward rules (ScherpelzLC07).
Our research highlights how prophecy variables can eliminate the need for backwards reasoning and separate treatment of forward and backward analyses — prophecy variables enable the unification of backward and forward analyses within a single forward reasoning framework.
Simulation relations, and techniques for proving that simulation relations exist, have been extensively explored in the context of establishing simulation relations between state machines (LynchV95; LynchDistributedAlgorithms). The developed theory includes a range of proof techniques and mechanisms, including forward and backward proof techniques with refinement mappings, abstraction functions, and abstraction relations. Prophecy variables were initially developed for the purpose of proving that implementations satisfy specifications via refinement mappings with forward simulations, specifically in the case when the specification makes a choice before the implementation (AbadiL91). The addition of prophecy variables to the framework of refinement mappings with history variables and forward simulation proofs enabled a completeness result for the ability to prove trace inclusions of implementations within specifications (AbadiL91). It is, of course, known that backward simulation is an alternative to forward simulation with prophecy variables (LynchV95). In general, there are a number of alternatives when choosing a formal framework for proving simulation properties, with the appropriate framework depending on pragmatic issues such as the convenience and conceptual difficulty of working with the concepts in the framework. In general, approaches that reason forward in time seem to be more attractive and intuitive than approaches that reason backward against time, as can be seen, for example, in pedagogical presentations of dataflow analyses, which invariably present forward analyses first, then backward analyses second as a kind of dual of forward analyses (appel2004modern; muchnick1997advanced; DragonBook; cooper2011engineering; kennedy2001optimizing; Aldrich2019Correctness; Aldrich2019Examples).
Prophecy variables have recently been applied for program verification in a Hoare program logic based on separation logic (JungLPRTDJ20). They have also been used as a mechanism to support reasoning about transformations between different program representations (LalR08; CookK11). Our purpose is different, specifically to use prophecy variables to enable forward reasoning and eliminate the need for backwards analyses and explicit intermediate representations in the context of a lightweight DSL implementation system that uses information about the future execution of the program to optimize the generated code.
5.2. Domain Specific Language Implementations
Many DSLs provide domain abstractions directly in the programming language to enhance developer productivity. Examples include visualization/graphics languages such as Tikz  (tikz), Vega/Vega-Lite (vega; vegalite), gnuplot (gnuplot), matplotlib (matplotlib), and Graphviz (graphviz); build system and build system generation languages such as make (make), cmake (cmake), and Ninja (ninja); and lexing and parsing languages like flex/lex bison/yacc (flex; yacc) and ANTLR (antlr). Because the main focus of these languages is increased developer productivity and not performance, implementations typically perform little to no forwards or backwards analysis to enable optimizations. Our research differs in that it uses prophecy variables for transformations and optimizations that require information about the future execution of the program.
DSLs such as GraphIt (graphit; g2; ugf), TACO (taco), Halide (halide), Diderot (diderot), Taichi (taichi), SQL (mysql; sqlite) and others target domains in which performance is a primary concern. Implementations of these DSLs therefore often contain a range of both forward and backward program analyses to enable the generation of optimized code. Implementations typically feature standard programming language implementation components such as parsers, intermediate representations, and forwards and backwards analyses that operate on these intermediate representations. Our research differs in that it uses the lightweight BuildIt approach, which eliminates many of these components. The combination of prophecy variables and reexecution to deliver correct prophecy variable values enables BuildIt to implement transformations and optimizations that require information about the future execution of the program using only forward execution, with no backwards analysis and no intermediate representation.
Partial evalution/multistage programming is an established language implementation technique with a long history of research (Futamura1971; JonesGomardSestoft1993; ConselDanvy1993; DaviesPfenning1996; TahaSheard1997; TahaSheard2000; CaretteKiselyovShan2009; RompfOdersky2010; Danvy1996; Taha2004). Examples of general purpose languages with multistage support include Template Haskell (sheard2002template), Lightweight Modular Staging in Scala (lms), Scheme (scheme), MetaML (TahaSheard2000), Meta OCaml, Mint (mint), Terra (terra; terra2) implemented in Lua, C macros, and C++ templates (vandevoorde2002c++). Multistage programming is often used to embed a DSL within a general purpose language. Examples include PyTorch (pytorch), Tensorflow (tensorflow), TVM (Chen2018TVM), AnyDSL (anydsl), Delite (delite), LMS (lms), BuilDSL (buildsl), JAX (jax2018github), SHiM (shim), and BREeze (tamara:meng-thesis:2023). Such language implementations often include sophisticated program analyses, either via standard programming language implementation techniques, or using history variables or their equivalent (the best of our knowledge, none of these systems use prophecy variables or their equivalent). Our research differs in that it 1) uses a lightweight approach with no intermediate representation and no program analysis of this (nonexistent) intermediate representation, 2) uses prophecy variables to enable transformations and optimizations that require information about the future execution of the program, and 3) uses repeated first stage executions to compute correct prophecy variable predictions, making it possible to implement transformations and optimizations that require information about the future execution of the program in a staged compilation system built on a standard programming language implementation that provides only forward program execution.
6. Conclusion
The staged BuildIt lightweight DSL implementation system delivers a combination of substantially reduced implementation effort and performance that is comparable to and often exceeds the performance of heavyweight standard DSL implementations. BuildIt is implemented as C++ data structures that encapsulate sophisticated program information extraction, transformation, and multioperation optimization techniques. This approach eliminates otherwise required program language implementation components such as parsers and intermediate representations. Instead of using forwards and backwards dataflow analysis to extract information, BuildIt uses history variables to record and propagate information about the past execution of the program to BuildIt first stage program transformation and optimization sites.
We introduce the combination of prophecy variables and repeated program execution, with the repeated program execution producing correct prophecy variable predictions. Our results illustrate how this combination extends the scope of the BuildIt implementation to include transformations and optimizations that require information about the future program execution while preserving the substantial engineering effort advantages of the basic BuildIt approach. By extending the scope of BuildIt to include languages whose efficient implementation requires these kinds of transformations and optimizations, the presented techniques promise to further eliminate barriers to rapid and efficient DSL implementation, enabling more developers to benefit from these effective application implementation platforms.
Supplemental Material for Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System
6.1. Einsum DSL Complete Implementation
This supplementary material provides the complete source code for the Einsum DSL implemented on top of BuildIt. This implementation uses a combination of prophecy variables and history variables to identify tensors to be moved between CPU and GPU. The implementation only has the BuildIt multi-stage library as a dependency and is compiled as regular C++ code. The source code below also includes the DSL program for matrix multiplication and a main function to generate code for the DSL.
1#include "builder/dyn_var.h"
2#include "builder/static_var.h"
3#include "builder/generics.h"
4#include "builder/nd_var.h"
5#include "blocks/extract_cuda.h"
6#include "builder/builder_context.h"
7#include "blocks/c_code_generator.h"
8#include "builder/signature_extract.h"
9
10
11namespace el {
12template <typename T>
13using prophecy_var = builder::nd_var<T>;
14using builder::with_name;
15
16namespace runtime {
17builder::dyn_var<void*(int)> malloc = with_name("runtime::typed_malloc");
18builder::dyn_var<void(void*)> free = with_name("runtime::typed_free");
19builder::dyn_var<void(void)> start_time = with_name("runtime::start_time");
20builder::dyn_var<void(void)> end_time = with_name("runtime::end_time");
21builder::dyn_var<void(void)> grid_sync = with_name("runtime::grid_sync");
22builder::dyn_var<void(void*, int)> cuda_malloc = with_name("runtime::cuda_malloc");
23builder::dyn_var<void(void*, int)> unified_malloc = with_name("runtime::unified_malloc");
24builder::dyn_var<void(void*, void*, int)> cudaMemcpyToDevice = with_name("runtime::cudaMemcpyToDevice");
25builder::dyn_var<void(void*, void*, int)> cudaMemcpyToHost = with_name("runtime::cudaMemcpyToHost");
26}
27
28static constexpr const int max_bid = 40;
29static constexpr const int max_tid = 512;
30
31enum device_type {
32 device_cpu,
33 device_gpu
34};
35builder::dyn_var<int> *current_bid = nullptr;
36builder::dyn_var<int> *current_tid = nullptr;
37device_type current_device = device_cpu;
38
39bool option_copy_all_tensors = false;
40bool option_use_unified_memory = false;
41
42struct index;
43struct tensor_base;
44template <typename T> struct tensor;
45template <typename T> struct tensor_access;
46template <typename T> struct tensor_term;
47
48struct index {
49 builder::dyn_var<int> *m_it;
50 // Index ranges are decided when they are used
51 int m_range = -1;
52};
53
54struct tensor_base {
55 static std::vector<tensor_base*> registered_tensors;
56
57 prophecy_var<builder::true_top> needs_gpu = builder::true_top::F;
58
59 builder::static_var<bool> gpu_written = false;
60 prophecy_var<builder::true_top> *gpu_read = nullptr;
61
62 tensor_base() {
63 registered_tensors.push_back(this);
64 }
65
66 virtual ~tensor_base();
67
68 virtual void move_to_gpu() = 0;
69 virtual void move_to_host() = 0;
70};
71
72// Type declarations
73template <typename T>
74struct tensor: public tensor_base {
75 std::vector<int> m_sizes;
76 int m_elemsize;
77
78 builder::dyn_var<T*> m_buffer;
79 builder::dyn_var<T*> m_gpu_buffer;
80
81 tensor(const std::vector<int> &sizes): m_sizes(sizes), m_elemsize(sizeof(T)) {
82 if (option_use_unified_memory) {
83 runtime::unified_malloc(&m_buffer, get_total_size() * m_elemsize);
84 } else {
85 runtime::malloc(&m_buffer, get_total_size() * m_elemsize);
86 if (option_copy_all_tensors needs_gpu.get()->value == builder::true_top::T) {
87 runtime::cuda_malloc(&m_gpu_buffer, get_total_size() * m_elemsize);
88 }
89 }
90 }
91
92 tensor(const std::vector<int> &sizes, const builder::dyn_var<T*>& buffer): m_sizes(sizes),
93 m_elemsize(sizeof(T)) {
94 // if user supplies buffer, it is users responsibility to use unified memory if necessary
95 if (!option_use_unified_memory) {
96 if (option_copy_all_tensors needs_gpu.get()->value == builder::true_top::T) {
97 runtime::cuda_malloc(&m_gpu_buffer, get_total_size() * m_elemsize);
98 }
99 }
100 }
101 ~tensor() {}
102
103 int get_total_size();
104
105 // Delete the copy constructors and assignments
106 tensor(const tensor&) = delete;
107 tensor(tensor&&) = delete;
108 tensor& operator=(const tensor&) = delete;
109 tensor& operator=(tensor&&) = delete;
110
111 // operations
112 tensor_access<T> operator[] (index& i);
113
114 void move_to_gpu() override;
115 void move_to_host() override;
116};
117
118template <typename T>
119struct tensor_access {
120 tensor<T>& m_tensor;
121 std::vector<index*> m_indices;
122
123 tensor_access(tensor<T>& t, index& i): m_tensor(t), m_indices({&i}) {}
124
125 void operator=(const tensor_term<T>&);
126 void operator+=(const tensor_term<T>&);
127 void operator*=(const tensor_term<T>&);
128
129 void operator=(const tensor_access<T>& o);
130
131 tensor_access<T> operator[] (index& i) const;
132
133 std::vector<index*> gather_indices(void) const;
134 void gather_ranges(void) const;
135
136 std::pair<std::vector<index*>, std::vector<index*>> initialize_indices(const tensor_term<T>&rhs);
137
138 enum reduction_op {red_assign, red_sum, red_product};
139 void induce_lhs_loops(const std::vector<index*> &lhs_indices, int ind,
140 enum reduction_op op, const std::vector<index*> &rhs_indices, const tensor_term<T>& rhs) const;
141 void induce_rhs_loops(const std::vector<index*> &rhs_indices, int ind, reduction_op op,
142 const tensor_term<T>& rhs, builder::dyn_var<T>* agg) const;
143
144 builder::dyn_var<int> get_flat_index(int index) const;
145
146};
147
148template <typename T>
149struct tensor_term {
150 enum mode {
151 mode_tensor,
152 mode_value,
153 mode_sum,
154 mode_product,
155 mode_index,
156 };
157 mode m_mode;
158 const tensor_access<T>* m_tensor;
159 // allow the m_value to be mutable because the type is set later
160 builder::dyn_var<T> m_value;
161 std::unique_ptr<const tensor_term<T>> m_e1;
162 std::unique_ptr<const tensor_term<T>> m_e2;
163 index* m_index;
164
165 // The 4 constructors, sum and product share a constructor
166 // The builder with_name trick is used to suppress the construction of m_value when not required
167 // We could also used defer_init
168 tensor_term(const tensor_access<T>& ta): m_mode(mode_tensor), m_tensor(&ta),
169 m_value(builder::with_name("__discard")) {}
170 tensor_term(tensor_term&& e1, tensor_term&& e2):
171 m_e1(std::unique_ptr<tensor_term<T>>(new tensor_term<T>(std::move(e1)))),
172 m_e2(std::unique_ptr<tensor_term<T>>(new tensor_term<T>(std::move(e2)))),
173 m_value(builder::with_name("__discard")) {
174 }
175 tensor_term(index &i): m_mode(mode_index), m_index(&i),
176 m_value(builder::with_name("__discard")) {}
177
178 // A catch all constructor for anything that is convertible to dyn_var
179 template <typename T2>
180 tensor_term (const T2& t): m_mode(mode_value), m_value(t) {}
181
182 std::vector<index*> gather_indices(void) const;
183 void gather_ranges(void) const;
184
185 builder::dyn_var<T> get_value() const;
186};
187
188// Type method definitions
189template <typename T>
190tensor_access<T> tensor<T>::operator[] (index &i) {
191 return tensor_access(*this, i);
192}
193template <typename T>
194int tensor<T>::get_total_size(void) {
195 int total = 1;
196 for (auto s: m_sizes) {
197 total *= s;
198 }
199 return total;
200}
201
202template <typename T>
203tensor_access<T> tensor_access<T>::operator[] (index& i) const {
204 tensor_access ret (*this);
205 ret.m_indices.push_back(&i);
206 return ret;
207}
208template <typename T>
209std::pair<std::vector<index*>, std::vector<index*>>
210tensor_access<T>::initialize_indices(const tensor_term<T>&rhs) {
211 // Initialize indices for acquiring range
212 std::vector<index*> lhs_indices = gather_indices();
213 std::vector<index*> rhs_indices = rhs.gather_indices();
214 for (auto a: lhs_indices) {
215 a->m_range = -1;
216 }
217 for (auto a: rhs_indices) {
218 a->m_range = -1;
219 }
220
221 gather_ranges();
222 rhs.gather_ranges();
223
224 // At this point make sure all indices used have ranges
225 for (auto a: lhs_indices) {
226 assert(a->m_range != -1);
227 }
228 for (auto a: rhs_indices) {
229 assert(a->m_range != -1);
230 }
231
232 std::vector<index*> reduce_indices;
233 for (auto a: rhs_indices) {
234 if (std::find(lhs_indices.begin(), lhs_indices.end(), a) == lhs_indices.end()) {
235 reduce_indices.push_back(a);
236 }
237 }
238 return std::make_pair(lhs_indices, reduce_indices);
239}
240
241template <typename T>
242builder::dyn_var<int> tensor_access<T>::get_flat_index(int ind) const {
243 if ((ind + 1) == m_indices.size()) return *(m_indices[ind]->m_it);
244 int size_after = 1;
245 for (unsigned i = ind + 1; i < m_tensor.m_sizes.size(); i++) {
246 size_after *= m_tensor.m_sizes[ind];
247 }
248 return (builder::cast)(size_after * *(m_indices[ind]->m_it) + get_flat_index(ind + 1));
249}
250
251template <typename T>
252void tensor_access<T>::induce_rhs_loops(const std::vector<index*> &rhs_indices, int ind, reduction_op op,
253 const tensor_term<T>& rhs, builder::dyn_var<T>* agg) const {
254 if (ind == rhs_indices.size()) {
255 if (op == red_product) {
256 *agg = *agg * rhs.get_value();
257 } else if (op == red_sum) {
258 *agg = *agg + rhs.get_value();
259 }
260 return;
261 }
262 for (builder::dyn_var<int> i = 0; i < rhs_indices[ind]->m_range; ++i) {
263 rhs_indices[ind]->m_it = i.addr();
264 induce_rhs_loops(rhs_indices, ind + 1, op, rhs, agg);
265 }
266 rhs_indices[ind]->m_it = nullptr;
267}
268
269template <typename T>
270void tensor_access<T>::induce_lhs_loops(const std::vector<index*> &lhs_indices, int ind, reduction_op op,
271 const std::vector<index*> &rhs_indices, const tensor_term<T>& rhs) const {
272 if (ind == lhs_indices.size()) {
273 builder::dyn_var<T*>* update_var;
274 if (current_device == device_gpu && !option_use_unified_memory) {
275 if (!option_copy_all_tensors)
276 m_tensor.needs_gpu.require_val(builder::true_top::T);
277 update_var = m_tensor.m_gpu_buffer.addr();
278 if (!option_copy_all_tensors && !option_use_unified_memory)
279 m_tensor.gpu_written = true;
280 } else {
281 update_var = m_tensor.m_buffer.addr();
282 }
283 if (op == red_assign) {
284 (*update_var)[get_flat_index(0)] = rhs.get_value();
285 } else if (op == red_product) {
286 builder::dyn_var<T> prod = 1;
287 induce_rhs_loops(rhs_indices, 0, op, rhs, prod.addr());
288 (*update_var)[get_flat_index(0)] = prod;
289 } else if (op == red_sum) {
290 builder::dyn_var<T> sum = 0;
291 induce_rhs_loops(rhs_indices, 0, op, rhs, sum.addr());
292 (*update_var)[get_flat_index(0)] = sum;
293 }
294 return;
295 }
296 // GPU Specializations
297 if (current_device == device_gpu) {
298 // If the lhs has only one loop, use the TID directly
299 if (lhs_indices.size() == 1) {
300 builder::dyn_var<int> thread = (*current_bid) * max_tid + (*current_tid);
301 for (builder::dyn_var<int> i = thread; i < lhs_indices[ind]->m_range; i += max_tid * max_bid) {
302 lhs_indices[ind]->m_it = i.addr();
303 induce_lhs_loops(lhs_indices, ind + 1, op, rhs_indices, rhs);
304 }
305 lhs_indices[ind]->m_it = nullptr;
306 runtime::grid_sync();
307 return;
308 } else {
309 // Special case for index 1 and 2
310 if (ind == 0) {
311 for (builder::dyn_var<int> i = *current_bid; i < lhs_indices[0]->m_range; i += max_bid) {
312 lhs_indices[0]->m_it = i.addr();
313 for (builder::dyn_var<int> j = *current_tid; j < lhs_indices[1]->m_range; j+= max_tid) {
314 lhs_indices[1]->m_it = j.addr();
315 induce_lhs_loops(lhs_indices, ind + 2, op, rhs_indices, rhs);
316 }
317 }
318 lhs_indices[0]->m_it = nullptr;
319 lhs_indices[1]->m_it = nullptr;
320 runtime::grid_sync();
321 return;
322 }
323 }
324 }
325
326 // to allow parallelization we have to use a copy
327 for (builder::dyn_var<int> i = 0; i < lhs_indices[ind]->m_range; ++i) {
328 lhs_indices[ind]->m_it = i.addr();
329 induce_lhs_loops(lhs_indices, ind + 1, op, rhs_indices, rhs);
330 }
331 lhs_indices[ind]->m_it = nullptr;
332}
333
334template <typename T>
335void tensor_access<T>::operator=(const tensor_term<T>& rhs) {
336 auto indices = initialize_indices(rhs);
337 std::vector<index*> lhs_indices = indices.first;
338 std::vector<index*> reduce_indices = indices.second;
339
340 // Specifically for assignment, reduce_indices should be empty
341 assert(reduce_indices.size() == 0);
342
343 induce_lhs_loops(lhs_indices, 0, red_assign, reduce_indices, rhs);
344}
345
346template <typename T>
347void tensor_access<T>::operator=(const tensor_access<T>& o) {
348 this->operator=((tensor_term<T>)o);
349}
350
351template <typename T>
352void tensor_access<T>::operator+=(const tensor_term<T>& rhs) {
353 auto indices = initialize_indices(rhs);
354 std::vector<index*> lhs_indices = indices.first;
355 std::vector<index*> reduce_indices = indices.second;
356 induce_lhs_loops(lhs_indices, 0, red_sum, reduce_indices, rhs);
357}
358
359template <typename T>
360void tensor_access<T>::operator*=(const tensor_term<T>& rhs) {
361 auto indices = initialize_indices(rhs);
362 std::vector<index*> lhs_indices = indices.first;
363 std::vector<index*> reduce_indices = indices.second;
364 induce_lhs_loops(lhs_indices, 0, red_product, reduce_indices, rhs);
365}
366
367template <typename T>
368std::vector<index*> tensor_access<T>::gather_indices(void) const {
369 // Make sure all indices are provided
370 assert(m_indices.size() == m_tensor.m_sizes.size());
371 // deduplicate indices and return
372 std::vector<index*> ret;
373 for (auto a: m_indices) {
374 if (std::find(ret.begin(), ret.end(), a) == ret.end()) {
375 ret.push_back(a);
376 }
377 }
378 return ret;
379}
380
381template <typename T>
382void tensor_access<T>::gather_ranges(void) const {
383 for (unsigned int i = 0; i < m_indices.size(); i++) {
384 auto ind = m_indices[i];
385 if (ind->m_range != -1) {
386 assert(ind->m_range == m_tensor.m_sizes[i]);
387 }
388 ind->m_range = m_tensor.m_sizes[i];
389 }
390}
391
392template <typename T>
393std::vector<index*> tensor_term<T>::gather_indices(void) const {
394 if (m_mode == mode_value) {
395 return {};
396 } else if (m_mode == mode_tensor) {
397 return m_tensor->gather_indices();
398 } else if (m_mode == mode_product m_mode == mode_sum) {
399 std::vector<index*> ret = m_e1->gather_indices();
400 std::vector<index*> other = m_e2->gather_indices();
401 for (auto a: other) {
402 if (std::find(ret.begin(), ret.end(), a) == ret.end()) {
403 ret.push_back(a);
404 }
405 }
406 return ret;
407 } else if (m_mode == mode_index) {
408 return {m_index};
409 }
410 return {};
411}
412
413template <typename T>
414void tensor_term<T>::gather_ranges(void) const {
415 if (m_mode == mode_tensor) {
416 m_tensor->gather_ranges();
417 } else if (m_mode == mode_product m_mode == mode_sum) {
418 m_e1->gather_ranges();
419 m_e2->gather_ranges();
420 }
421 // index doesn’t gather ranges
422}
423
424template <typename T>
425builder::dyn_var<T> tensor_term<T>::get_value(void) const {
426 if (m_mode == mode_value) {
427 return m_value;
428 } else if (m_mode == mode_tensor) {
429 if (current_device == device_gpu && !option_use_unified_memory) {
430 if (!option_copy_all_tensors)
431 m_tensor->m_tensor.needs_gpu.require_val(builder::true_top::T);
432 if (!option_copy_all_tensors && !option_use_unified_memory)
433 m_tensor->m_tensor.gpu_read->require_val(builder::true_top::T);
434 return m_tensor->m_tensor.m_gpu_buffer[m_tensor->get_flat_index(0)];
435 } else {
436 return m_tensor->m_tensor.m_buffer[m_tensor->get_flat_index(0)];
437 }
438 } else if (m_mode == mode_product) {
439 return m_e1->get_value() * m_e2->get_value();
440 } else if (m_mode == mode_sum) {
441 return m_e1->get_value() + m_e2->get_value();
442 } else if (m_mode == mode_index) {
443 return *(m_index->m_it);
444 }
445 return 0;
446}
447template <typename T>
448void tensor<T>::move_to_gpu() {
449 runtime::cudaMemcpyToDevice(m_gpu_buffer, m_buffer, get_total_size() * m_elemsize);
450}
451template <typename T>
452void tensor<T>::move_to_host() {
453 runtime::cudaMemcpyToHost(m_buffer, m_gpu_buffer, get_total_size() * m_elemsize);
454}
455
456// utility SFINAE to check if a type is
457// one of the types that can appear in expressions
458
459template <typename T>
460struct is_tensor_expr: std::false_type {};
461template <typename T>
462struct is_tensor_expr<tensor_term<T>>: std::true_type {
463 using wrapped_type = T;
464};
465template <typename T>
466struct is_tensor_expr<tensor_access<T>>: std::true_type {
467 using wrapped_type = T;
468};
469
470template <typename T1, typename T2, typename V=void>
471struct is_valid_arg_types {};
472
473template <typename T1, typename T2>
474struct is_valid_arg_types<T1, T2, typename std::enable_if<is_tensor_expr<T1>::value>::type> {
475 using type = tensor_term<typename is_tensor_expr<T1>::wrapped_type>;
476};
477template <typename T1, typename T2>
478struct is_valid_arg_types<T1, T2, typename std::enable_if<!is_tensor_expr<T1>::value &&
479 is_tensor_expr<T2>::value>::type> {
480 using type = tensor_term<typename is_tensor_expr<T2>::wrapped_type>;
481};
482
483
484template <typename T1, typename T2>
485auto operator + (T1&& e1, T2&& e2) -> typename is_valid_arg_types<T1, T2>::type {
486 using $t$ = typename is_valid_arg_types<T1, T2>::type;
487 $t$ ret(e1, e2);
488 ret.m_mode = $t$::mode_sum;
489 return ret;
490}
491template <typename T1, typename T2>
492auto operator * (T1&& e1, T2&& e2) -> typename is_valid_arg_types<T1, T2>::type {
493 using $t$ = typename is_valid_arg_types<T1, T2>::type;
494 $t$ ret(std::move(e1), std::move(e2));
495 ret.m_mode = $t$::mode_product;
496 return ret;
497}
498
499template <typename RetType, typename ArgTuple>
500struct wrap_helper;
501
502template <typename RetType, typename...Args>
503struct wrap_helper<RetType, std::tuple<Args...>> {
504 static std::function<RetType(Args...)> wrap (std::function<RetType(Args...)> f) {
505 return [=](Args...args) -> RetType {
506 current_device = device_cpu;
507 return f(args...);
508 };
509 }
510};
511
512template <typename...Args>
513struct wrap_helper<void, std::tuple<Args...>> {
514 static std::function<void(Args...)> wrap (std::function<void(Args...)> f) {
515 return [=](Args...args) -> void {
516 current_device = device_cpu;
517 f(args...);
518 };
519 }
520};
521
522
523
524void run_on_gpu(std::function<void(void)> func);
525void run_on_gpu(std::function<void(void)> func) {
526 current_device = device_gpu;
527
528 // Before launching the kernel, non-deterministically guess which
529 // tensors are required on the gpu
530
531 for (builder::static_var<unsigned> i = 0; i < tensor_base::registered_tensors.size(); i++) {
532 auto t = tensor_base::registered_tensors[i];
533 if (!option_use_unified_memory) {
534 if (option_copy_all_tensors) {
535 t->move_to_gpu();
536 } else {
537
538 if (t->gpu_read) delete t->gpu_read;
539 t->gpu_written = false;
540 t->gpu_read = new prophecy_var<builder::true_top>(builder::true_top::F);
541
542
543 if (t->gpu_read->get()->value == builder::true_top::T) {
544 t->move_to_gpu();
545 }
546 }
547 }
548 }
549
550 builder::annotate(CUDA_KERNEL_COOP);
551 for (builder::dyn_var<int> bid = 0; bid < max_bid; ++bid) {
552 for (builder::dyn_var<int> tid = 0; tid < max_tid; ++tid) {
553 current_bid = bid.addr();
554 current_tid = tid.addr();
555 func();
556 }
557 }
558 current_bid = nullptr;
559 current_tid = nullptr;
560 current_device = device_cpu;
561
562 // After finishing copy back tensors that have been written and delete the nd vars
563
564 for (builder::static_var<unsigned> i = 0; i < tensor_base::registered_tensors.size(); i++) {
565 auto t = tensor_base::registered_tensors[i];
566
567 if (!option_use_unified_memory) {
568 if (option_copy_all_tensors) {
569 t->move_to_host();
570 } else {
571 if (t->gpu_written) {
572 t->move_to_host();
573 }
574
575 delete t->gpu_read;
576 t->gpu_written = false;
577 t->gpu_read = nullptr;
578 }
579 }
580 }
581}
582
583template <typename T, typename...Args>
584void generate_code(std::ostream& oss, const T& func, std::string fname, Args&&...args) {
585 builder::builder_context ctx;
586 ctx.run_rce = true;
587 using ArgTypes = typename builder::f_type_helper<T>::arg_types;
588 using ReturnType = typename builder::f_type_helper<T>::ret_type;
589 auto wrapped = wrap_helper<ReturnType, ArgTypes>::wrap(func);
590 auto ast = ctx.extract_function_ast(wrapped, fname, std::forward<Args>(args)...);
591
592 auto new_decls = block::extract_cuda_from(block::to<block::func_decl>(ast)->body);
593
594 oss << "#include \"el_runtime.h\"" << std::endl;
595 for (auto a: new_decls)
596 block::c_code_generator::generate_code(a, oss, 0);
597 block::c_code_generator::generate_code(ast, oss, 0);
598}
599
600std::vector<tensor_base*> tensor_base::registered_tensors;
601
602tensor_base::~tensor_base() {
603 registered_tensors.erase(std::remove(registered_tensors.begin(), registered_tensors.end(), this),
604 registered_tensors.end());
605 if (gpu_read) delete gpu_read;
606}
607
608}
609
610void benchmark (builder::dyn_var<float*> _x, builder::dyn_var<float*> _y,
611 builder::dyn_var<float*> _z, int M, int N, int O) {
612 el::index i, j, k;
613
614 el::tensor<float> x({M, N});
615 el::tensor<float> x_b({M, N}, _x);
616 el::tensor<float> y({N, O});
617 el::tensor<float> y_b({N, O}, _y);
618 el::tensor<float> z({M, O});
619
620
621 // temporary to read data into
622 el::tensor<float> z_b({M, O}, _z);
623
624
625 for (builder::dyn_var<int> iter = 0; iter < 10; iter++) {
626
627 x[i][j] = x_b[i][j];
628 y[i][j] = y_b[i][j];
629
630 el::runtime::start_time();
631
632 el::run_on_gpu([&]() {
633 z[i][j] += x[i][k] * y[k][j];
634 });
635
636 z_b[i][j] = z[i][j];
637
638 el::runtime::end_time();
639 }
640
641}
642
643
644int main(int argc, char* argv[]) {
645 el::generate_code(std::cout, benchmark, "benchmark", 1024, 1024, 1024);
646 el::option_copy_all_tensors = true;
647 el::generate_code(std::cout, benchmark, "benchmark_copy_all", 1024, 1024, 1024);
648 el::option_use_unified_memory = true;
649 el::generate_code(std::cout, benchmark, "benchmark_unified", 1024, 1024, 1024);
650 return 0;
651}
6.2. Neural Network Fusion DSL Complete Implementation
This supplementary material provides the complete source code for the mini neural network DSL implemented on top of BuildIt. This implementation uses a combination of prophecy variables and history variables to automatically fuse nonlinearities into complex operations like convolutions. The implementation only has the BuildIt multi-stage library as a dependency and is compiled as regular C++ code. The source code below also includes the DSL program for simple test where the operations cannot be fused due to ReLU with different thresholds on different paths and an example where it can be used.
1#include "builder/dyn_var.h"
2#include "builder/static_var.h"
3#include "builder/builder_context.h"
4#include "blocks/c_code_generator.h"
5#include "builder/nd_var.h"
6
7namespace ml {
8template <typename T>
9using prophecy_var = builder::nd_var<T>;
10
11namespace runtime {
12builder::dyn_var<void*(int)> malloc = builder::with_name("malloc");
13builder::dyn_var<void(void*)> free = builder::with_name("free");
14builder::dyn_var<void(void*, void*, int)> memcpy = builder::with_name("memcpy");
15builder::dyn_var<void(void)> start_time = builder::with_name("start_time");
16builder::dyn_var<void(void)> end_time = builder::with_name("end_time");
17builder::dyn_var<void(void)> printf = builder::with_name("printf");
18builder::dyn_var<void(float*)> consume_tensor = builder::with_name("consume_tensor");
19}
20// A simple true at top boolean nd_var wrappable type
21class false_top: public builder::nd_var_base {
22public:
23 typedef enum {
24 F = 2,
25 T = 1,
26 Unspecified = 0,
27 } level_t;
28
29 typedef struct {
30 level_t level;
31 float threshold;
32 } value_type;
33 value_type value;
34
35 false_top(value_type def): value(def) {}
36 false_top(): value({Unspecified, 0.0}) {}
37
38 bool check(value_type e) {
39 if (value.level > e.level)
40 return true;
41 if (value.level < e.level)
42 return false;
43 if (value.level == T)
44 return abs(value.threshold - e.threshold) < 0.001;
45 return true;
46 }
47 void merge(value_type e) {
48 // extra sanity check
49 if (check(e)) return;
50 if (value.level == T && e.level == T) {
51 value.level = F;
52 return;
53 }
54 value = e;
55 }
56};
57
58struct tensor {
59 int m_size;
60 builder::dyn_var<float*> m_array;
61
62 builder::static_var<bool> is_last_convolution = false;
63 prophecy_var<false_top>* is_next_relu = nullptr;
64
65
66 tensor(builder::dyn_var<float*> array, const int size): m_size(size), m_array(array) {}
67 tensor(const int size): m_size(size), m_array(runtime::malloc(size * sizeof(float))) {}
68 tensor(tensor& other): m_size(other.m_size), m_array(runtime::malloc(m_size * sizeof(float))) {
69 runtime::memcpy(m_array, other.m_array, m_size * sizeof(float));
70 }
71 // move constructor canibalizes the other tensor
72 tensor(tensor&& other): m_size(other.m_size), m_array(other.m_array) {}
73
74 void operator=(tensor& other) {
75 if (m_size != other.m_size) {
76 runtime::free(m_array);
77 m_size = other.m_size;
78 m_array = runtime::malloc(m_size * sizeof(float));
79 }
80 runtime::memcpy(m_array, other.m_array, m_size * sizeof(float));
81 }
82 void operator=(tensor&& other) {
83 runtime::free(m_array);
84 m_size = other.m_size;
85 m_array = other.m_array;
86 }
87};
88
89tensor convolve(tensor &input, tensor &filter);
90tensor convolve(tensor &input, tensor &filter) {
91 tensor output (input.m_size);
92
93 output.is_last_convolution = true;
94 output.is_next_relu = new prophecy_var<false_top>();
95
96 // simple convolution that follows wrap around behavior
97 for (builder::dyn_var<int> i = 0; i < input.m_size; i++) {
98 builder::dyn_var<float> sum = 0;
99 for (builder::dyn_var<int> j = 0; j < filter.m_size; j++) {
100 sum += input.m_array[(i + j) % input.m_size] * filter.m_array[j];
101 }
102 if (output.is_next_relu->get()->value.level == false_top::T) {
103 if (sum < output.is_next_relu->get()->value.threshold) {
104 sum = 0;
105 }
106 }
107 output.m_array[i] = sum;
108 }
109 return output;
110}
111tensor relu(tensor &input, const float threshold);
112tensor relu(tensor &input, const float threshold) {
113 if (input.is_last_convolution) {
114 input.is_last_convolution = false;
115 input.is_next_relu->require_val({false_top::T, threshold});
116 if (input.is_next_relu->get()->value.level == false_top::T)
117 return std::move(input);
118 }
119 for (builder::dyn_var<int> i = 0; i < input.m_size; i++) {
120 if (input.m_array[i] >= threshold) {
121 input.m_array[i] = input.m_array[i];
122 } else {
123 input.m_array[i] = 0;
124 }
125 }
126 return std::move(input);
127}
128}
129
130static builder::dyn_var<float*> benchmark(builder::dyn_var<float*> _input, const int size,
131 builder::dyn_var<float*> _weight, const int weight_size, builder::dyn_var<bool> small_t) {
132 ml::tensor input(_input, size);
133 ml::tensor weight(_weight, weight_size);
134
[…]
136 for (builder::dyn_var<int> i = 0; i < 10; i++) {
137 ml::runtime::start_time();
138 ml::tensor output_conv = ml::convolve(input, weight);
139 if (small_t) {
140 ml::tensor output = ml::relu(output_conv, 2.0);
141 ml::runtime::end_time();
142 ml::runtime::consume_tensor(output.m_array);
143 } else {
144 // Different threshold on a different branch
145 ml::tensor output = ml::relu(output_conv, 4.0);
146 ml::runtime::end_time();
147 ml::runtime::consume_tensor(output.m_array);
148 }
149 }
150
151
[…]
153 for (builder::dyn_var<int> i = 0; i < 10; i++) {
154 ml::runtime::start_time();
155 ml::tensor output_conv = ml::convolve(input, weight);
156 ml::tensor output = ml::relu(output_conv, 1.56);
157 ml::runtime::end_time();
158 ml::runtime::consume_tensor(output.m_array);
159 }
160 return 0;
161}
162
163
164int main(int argc, char* argv[]) {
165
166 builder::builder_context ctx;
167 ctx.run_rce = true;
168
169 auto ast = ctx.extract_function_ast(benchmark, "benchmark", 10240, 21);
170 std::cout << "#include \"ml_runtime.h\"" <<std::endl;
171 block::c_code_generator::generate_code(ast, std::cout, 0);
172
173
174 return 0;
175}

### v2

@@
-Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System
+Backwards Data-Flow Analysis using Prophecy Variables in the BuildIt System
 Abstract.
@@
 BuildIt Standard DSL BuildIt Geometric Mean, Max, Min
-GraphIt (g2) / GraphIt (BuildIt)(buildsl) 23,783 2,021 3%, 33%, -8%
-COLA (shim) / SHIM (BuildIt) (shim) 7,500 5,700 1.4%, 32%, -30%
-PCRE2(pcre2) / BREze (BuildIt) (tamara:meng-thesis:2023) 133,995 1,564 800%, 9100%, -19%
-RE2(re2) / BREze (BuildIt) (tamara:meng-thesis:2023) 26,587 1,564 37%, 600%, -129%
-Hyperscan(hyperscan) / BREze (BuildIt) (tamara:meng-thesis:2023) 187,033 1,564 -127%, 2000%, -1700%
+GraphIt (Brahmakshatriya et al., 2021b) / GraphIt (BuildIt)(Brahmakshatriya and Amarasinghe, 2022) 23,783 2,021 3%, 33%, -8%
+COLA (Ray, 2023) / SHIM (BuildIt) (Ray, 2023) 7,500 5,700 1.4%, 32%, -30%
+PCRE2(Project, 2025) / BREze (BuildIt) (Mitrovska, 2023) 133,995 1,564 800%, 9100%, -19%
+RE2(Cox, 2009) / BREze (BuildIt) (Mitrovska, 2023) 26,587 1,564 37%, 600%, -129%
+Hyperscan(Wang et al., 2019) / BREze (BuildIt) (Mitrovska, 2023) 187,033 1,564 -127%, 2000%, -1700%
 Table 1.
@@
 These numbers show that BuildIt delivers performance that is at least comparable to and often significantly better than the performance provided by standard DSL implementations.
-Component BuildIt(buildsl) GraphIt(g2)
+Component BuildIt(Brahmakshatriya and Amarasinghe, 2022) GraphIt(Brahmakshatriya et al., 2021b)
 Frontend - 9,593
@@
 1.1. BuildIt
-In response to this situation, we developed BuildIt (buildsl; ajay-thesis), a system that leverages modern programming language features such as overloading to implement DSLs.
+In response to this situation, we developed BuildIt (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025), a system that leverages modern programming language features such as overloading to implement DSLs.
 Operating completely within a modern general purpose language, BuildIt implements compilation and optimization techniques that operate across the entire domain specific multioperation computation.
@@
 These numbers show that BuildIt implementations can require up to 100 times fewer lines of code than standard implementations of the same DSL, a reduction substantial enough to transform the viability of many DSLs.
-And the performance results in Table 2 show that, in comparison with standard DSL implementations, BuildIt implementations typically exhibit comparable or better performance (buildsl; ajay-thesis).
+And the performance results in Table 2 show that, in comparison with standard DSL implementations, BuildIt implementations typically exhibit comparable or better performance (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025).
 1.2. Program Analysis in BuildIt
@@
 These take place during the BuildIt first phase and typically target the entire multioperation computation.
-For analyses that propagate information forward with the flow of control, BuildIt uses history variables (implemented as data structure fields in the underlying implementation language) to carry the information through the computation (OwickiG76; AbadiL91; rinard2020dataflow).
+For analyses that propagate information forward with the flow of control, BuildIt uses history variables (implemented as data structure fields in the underlying implementation language) to carry the information through the computation (Owicki and Gries, 1976; Abadi and Lamport, 1991; Rinard and Gadient, 2020).
 These variables are assigned as the BuildIt first stage executes and carry the information through the first stage execution to be later accessed (again during the first stage) to optimize the generated second stage code.
@@
 Some optimizations, however, require information about the future program execution (history variables record information about the past, not future, execution).
-One standard way to obtain this information is to build an internal intermediate program representation, then use a backwards analysis to propagate information against the flow of control to reach the point where it is applied to optimize the generated code (DragonBook; appel2004modern; muchnick1997advanced; cooper2011engineering; kennedy2001optimizing).
+One standard way to obtain this information is to build an internal intermediate program representation, then use a backwards analysis to propagate information against the flow of control to reach the point where it is applied to optimize the generated code (Aho et al., 2006; Appel, 2004; Muchnick and others, 1997; Cooper and Torczon, 2011; Kennedy and Allen, 2001).
 A drawback of attempting to apply this approach in the BuildIt context is the need to construct an appropriate program representation to enable the backwards analysis, an endeavor that would move the BuildIt implementation closer to standard DSLimplementations, increase the BuildIt implementation effort, and erode the implementation effort advantage that BuildIt currently exhibits over standard implementations of DSLs.
 We instead present a novel use of prophecy variables to obtain information about the future program execution.
-Prophecy variables were initially developed to prove forward simulation relations in state machine refinement proofs that involve reasoning about future program behavior (AbadiL91).
+Prophecy variables were initially developed to prove forward simulation relations in state machine refinement proofs that involve reasoning about future program behavior (Abadi and Lamport, 1991).
 The basic idea behind prophecy variables in the BuildIt context is to use prophecy variable predictions to predict information about the future execution, then, when the information becomes available later in the execution, use prophecy variable preconditions to check that the prophecy variable correctly predicts the information.
-In the refinement proof context, prophecy variable predictions are typically nondeterministic, with angelic nondeterminism enabling the refinement proof to go through if any of the nondeterministic predictions was correct (AbadiL91).
+In the refinement proof context, prophecy variable predictions are typically nondeterministic, with angelic nondeterminism enabling the refinement proof to go through if any of the nondeterministic predictions was correct (Abadi and Lamport, 1991).
 In the BuildIt context, we instead use repeated program executions to discover correct prophecy variable predictions.
@@
 2. Example
-Figure 1 presents example BuildIt code for tensor computations using the Einstein summation notation (einstein_foundation_1916) (einsum notation) written in C++ (einsum-buildit).
+Figure 1 presents example BuildIt code for tensor computations using the Einstein summation notation (Einstein, 1916) (einsum notation) written in C++ (BuildIt-lang, 2022).
 The example uses prophecy variables to predict tensors that may be accessed by a computation performed on a GPU.
@@
 3.1. Overview
-We formalize prophecy variables and the prophecy variable analysis using a core programming language inspired by Glynn Winskell’s imperative language IMP (Winskel93).
-Notable differences include the introduction of labels for all commands and the use of variables $V$ instead of locations ($\mathbf{Loc}$(rinard2020dataflow).
+We formalize prophecy variables and the prophecy variable analysis using a core programming language inspired by Glynn Winskell’s imperative language IMP (Winskel, 1993).
+Notable differences include the introduction of labels for all commands and the use of variables $V$ instead of locations ($\mathbf{Loc}$(Rinard and Gadient, 2020).
 Specifically, the formalization works with programs $P$ that contain labeled commands of the form $l:c\in P$, where $l,g\in L$, $c\in C$.
@@
 The progress property requires the analysis to produce correct prophecy variable predictions about all possible future executions (in the sense that the analysis results satisfy the prophecy variable preconditions that check for incorrect predictions).
-We note that if Preservation (Definition 3.1) and Progress (Definition 3.2) both hold, then the relation $\sim$ defined by $\left\langle l,\sigma\right\rangle\sim\left\langle l,\sigma,\beta_{\bullet l}\right\rangle$ is a bisimulation relation (Milner89; Benthem96) between standard and extended program configurations (rinard2020dataflow).
+We note that if Preservation (Definition 3.1) and Progress (Definition 3.2) both hold, then the relation $\sim$ defined by $\left\langle l,\sigma\right\rangle\sim\left\langle l,\sigma,\beta_{\bullet l}\right\rangle$ is a bisimulation relation (Milner, 1989; van Benthem, 1996) between standard and extended program configurations (Rinard and Gadient, 2020).
 3.2. Core Programming Language
@@
 3.5. Prophecy Variable Analysis via Repeated Program Execution
-It is possible to use a backwards dataflow analyses to obtain prophecy variable values that correctly predict future program executions (rinard2020dataflow).
+It is possible to use a backwards dataflow analyses to obtain prophecy variable values that correctly predict future program executions (Rinard and Gadient, 2020).
 These analyses, however, work with an explicit program representation to propagate information backwards against the flow of control.
@@
 4.1. BuildIt Overview
-BuildIt is a lightweight DSL implementation system (ajay-thesis).
+BuildIt is a lightweight DSL implementation system (Brahmakshatriya, 2025).
 BuildIt programs execute two stages.
@@
 Together, static_var<T> and prophecy<T> variables can be used to collect information about both the previous and future execution of the program, then transfer that information to relevant transformation and optimization sites, eliminating the need for traditional program analysis algorithms that operate on intermediate program representations.
-BuildIt also uses program reexecution to explore and generate control flow that involves dyn_var<T> expressions (buildsl; ajay-thesis).
+BuildIt also uses program reexecution to explore and generate control flow that involves dyn_var<T> expressions (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025).
 We repurpose this program reexecution capability to implement the prophecy variable analysis formalized in Section 3.
 4.3. Tensor Data Movement for GPU Kernels
-This benchmark involves a BuildIt DSL that can evaluate arbitrary multidimensional tensor operations specified as einsum expressions (einsum-buildit).
+This benchmark involves a BuildIt DSL that can evaluate arbitrary multidimensional tensor operations specified as einsum expressions (BuildIt-lang, 2022).
 In this DSL the the dimensions and sizes of each tensor are first stage values.
@@
 We present performance results for two computations: 1) matrix-matrix multiply (six tensors, three of which are ever resident on the GPU) and 2) matrix-vector multiply (two matrices and four vectors, with one matrix and one vector ever resident on the GPU).
-We consider two alternative GPU tensor movement implementations: 1) moving all tensors to the GPU before executing the GPU kernel (we acknowledge that this strategy becomes infeasible for applications that manipulate large numbers of tensors) and 2) using a runtime data movement strategy ( CUDA unified memory, which uses a page fault based mechanism to move data between CPU and GPU (Harris2017UnifiedMemory)).
+We consider two alternative GPU tensor movement implementations: 1) moving all tensors to the GPU before executing the GPU kernel (we acknowledge that this strategy becomes infeasible for applications that manipulate large numbers of tensors) and 2) using a runtime data movement strategy ( CUDA unified memory, which uses a page fault based mechanism to move data between CPU and GPU (Harris, 2017)).
 Table 8 presents the minimum execution time in microseconds for GPU matrix-matrix multiplication running with the three strategies and four matrix sizes across ten runs.
@@
 We report minimum executions times in microseconds across 10 runs.
-Table 9 presents the number of first stage program executions required to 1) fully explore all possible dynamic control flow paths (this is a baseline BuildIt technique for ensuring correct second stage code generation (buildit; ajay-thesis)) and 2) obtain correct prophecy variable values.
+Table 9 presents the number of first stage program executions required to 1) fully explore all possible dynamic control flow paths (this is a baseline BuildIt technique for ensuring correct second stage code generation (Brahmakshatriya and Amarasinghe, 2021; Brahmakshatriya, 2025)) and 2) obtain correct prophecy variable values.
 It also presents the first stage execution times.
@@
 Number of first stage program executions and first stage execution times for Matrix Matrix Multiply and Matrix Vector Multiply.
-We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
+We report both dynamic control flow exploration executions (Brahmakshatriya, 2025) and executions triggered by prophecy variable precondition violations.
 4.4. Neural Network Operation Fusion
@@
 Number of first stage program executions and first stage execution times.
-We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
+We report both dynamic control flow exploration executions (Brahmakshatriya, 2025) and executions triggered by prophecy variable precondition violations.
 We report execution times for the benchmark program in Figure 10 with the convolution/ReLU fusion optimization disabled and enabled.
@@
 Many of the concepts that appear in simulation relation proofs for state machines also appear in the program verification, dataflow analysis, and abstract interpretation literature.
-For example, history variables were first introduced in the program verification literature (OwickiG76), abstraction functions, originally introduced in the program verification literature (Hoare72), can be seen as a form of refinement mappings, and program analyses can be seen as establishing a simulation relation between an abstract interpretation of the program (which plays the role of the specification) and concrete executions of the program (which play the role of the implementation) (CousotC77; CousotC92).
-It is also known that, in this context, backward or reverse simulation relations can be used to establish the correspondence between backward analyses (which extract information about the future execution) and program executions (CousotC92; SchmidtS98).
-A recent (unpublished in any peer reviewed venue) manuscript introduces prophecy variables to enable forward reasoning about program analysis properties that involve the future execution of the program (rinard2020dataflow).
-To the best of our knowledge, this manuscript is the first to introduce prophecy variables for program analysis (here we contrast with the use of prophecy variables for program verification (JungLPRTDJ20; Vafeiadis08; ZhangFFSL12) as well as the traditional use of prophecy variables for proving forward simulation relations between state machines (AbadiL91)).
+For example, history variables were first introduced in the program verification literature (Owicki and Gries, 1976), abstraction functions, originally introduced in the program verification literature (Hoare, 1972), can be seen as a form of refinement mappings, and program analyses can be seen as establishing a simulation relation between an abstract interpretation of the program (which plays the role of the specification) and concrete executions of the program (which play the role of the implementation) (Cousot and Cousot, 1977, 1992).
+It is also known that, in this context, backward or reverse simulation relations can be used to establish the correspondence between backward analyses (which extract information about the future execution) and program executions (Cousot and Cousot, 1992; Schmidt and Steffen, 1998).
+A recent (unpublished in any peer reviewed venue) manuscript introduces prophecy variables to enable forward reasoning about program analysis properties that involve the future execution of the program (Rinard and Gadient, 2020).
+To the best of our knowledge, this manuscript is the first to introduce prophecy variables for program analysis (here we contrast with the use of prophecy variables for program verification (Jung et al., 2020; Vafeiadis, 2008; Zhang et al., 2012) as well as the traditional use of prophecy variables for proving forward simulation relations between state machines (Abadi and Lamport, 1991)).
 The manuscript uses prophecy variables to specify two program analysis problems, live variables and very busy expressions, that require information about the future execution of the program and uses a backwards dataflow analysis to obtain correct prophecy variable predictions.
-While we build on the formalism presented in this manuscript, our research differs in that 1) we use repeated forward program executions, not backwards dataflow analysis, to find correct prophecy variable predictions, 2) we implement prophecy variables in the BuildIt staged DSL implementation system ( (rinard2020dataflow) presents no implementation), 3) we show how prophecy variables enable the BuildIt first stage, which uses only standard forward program execution, to obtain and exploit information about the future execution of the program, and 4) we report results from BuildIt prophecy variable implementations.
+While we build on the formalism presented in this manuscript, our research differs in that 1) we use repeated forward program executions, not backwards dataflow analysis, to find correct prophecy variable predictions, 2) we implement prophecy variables in the BuildIt staged DSL implementation system ( (Rinard and Gadient, 2020) presents no implementation), 3) we show how prophecy variables enable the BuildIt first stage, which uses only standard forward program execution, to obtain and exploit information about the future execution of the program, and 4) we report results from BuildIt prophecy variable implementations.
 The overall result is the extension of the scope of the lightweight BuildIt implementation approach to include optimizations (such as those presented in Section 4) that require information about the future execution of the program.
-Cobalt enables compiler developers to specify a range of dataflow optimizations (such as constant propagation and partial dead assignment elimination) (LernerMC03).
+Cobalt enables compiler developers to specify a range of dataflow optimizations (such as constant propagation and partial dead assignment elimination) (Lerner et al., 2003).
 Each optimization is specified by a transformation pattern whose guard specifies a condition over sequences of actions in paths in the program representation that must hold for the transformation to be legal.
 Cobalt has separate constructs for specifying forward and backward optimizations — forward guards reason about forward properties, backward guards reason about backward properties.
-Rhodium was developed to eliminate the complexity of using temporal logic formulas and implicit dataflow facts to reason about paths in control-flow graphs (LernerMRC05).
+Rhodium was developed to eliminate the complexity of using temporal logic formulas and implicit dataflow facts to reason about paths in control-flow graphs (Lerner et al., 2005).
 Rhodium uses de facto abstraction functions (expressed as predicates over concrete program states) to establish the connection between concrete program states and dataflow facts and state extensions (a form of instrumented semantics) to support analyses that extract information about the past execution of the program not present in standard concrete program states.
-Like Cobalt, Rhodium has separate support for forward and backward analyses; subsequent work on automatically inferring correct propagation rules supports only forward rules (ScherpelzLC07).
+Like Cobalt, Rhodium has separate support for forward and backward analyses; subsequent work on automatically inferring correct propagation rules supports only forward rules (Scherpelz et al., 2007).
 Our research highlights how prophecy variables can eliminate the need for backwards reasoning and separate treatment of forward and backward analyses — prophecy variables enable the unification of backward and forward analyses within a single forward reasoning framework.
-Simulation relations, and techniques for proving that simulation relations exist, have been extensively explored in the context of establishing simulation relations between state machines (LynchV95; LynchDistributedAlgorithms).
+Simulation relations, and techniques for proving that simulation relations exist, have been extensively explored in the context of establishing simulation relations between state machines (Lynch and Vaandrager, 1995; Lynch, 1996).
 The developed theory includes a range of proof techniques and mechanisms, including forward and backward proof techniques with refinement mappings, abstraction functions, and abstraction relations.
-Prophecy variables were initially developed for the purpose of proving that implementations satisfy specifications via refinement mappings with forward simulations, specifically in the case when the specification makes a choice before the implementation (AbadiL91).
-The addition of prophecy variables to the framework of refinement mappings with history variables and forward simulation proofs enabled a completeness result for the ability to prove trace inclusions of implementations within specifications (AbadiL91).
-It is, of course, known that backward simulation is an alternative to forward simulation with prophecy variables (LynchV95).
+Prophecy variables were initially developed for the purpose of proving that implementations satisfy specifications via refinement mappings with forward simulations, specifically in the case when the specification makes a choice before the implementation (Abadi and Lamport, 1991).
+The addition of prophecy variables to the framework of refinement mappings with history variables and forward simulation proofs enabled a completeness result for the ability to prove trace inclusions of implementations within specifications (Abadi and Lamport, 1991).
+It is, of course, known that backward simulation is an alternative to forward simulation with prophecy variables (Lynch and Vaandrager, 1995).
 In general, there are a number of alternatives when choosing a formal framework for proving simulation properties, with the appropriate framework depending on pragmatic issues such as the convenience and conceptual difficulty of working with the concepts in the framework.
-In general, approaches that reason forward in time seem to be more attractive and intuitive than approaches that reason backward against time, as can be seen, for example, in pedagogical presentations of dataflow analyses, which invariably present forward analyses first, then backward analyses second as a kind of dual of forward analyses (appel2004modern; muchnick1997advanced; DragonBook; cooper2011engineering; kennedy2001optimizing; Aldrich2019Correctness; Aldrich2019Examples).
-Prophecy variables have recently been applied for program verification in a Hoare program logic based on separation logic (JungLPRTDJ20).
-They have also been used as a mechanism to support reasoning about transformations between different program representations (LalR08; CookK11).
+In general, approaches that reason forward in time seem to be more attractive and intuitive than approaches that reason backward against time, as can be seen, for example, in pedagogical presentations of dataflow analyses, which invariably present forward analyses first, then backward analyses second as a kind of dual of forward analyses (Appel, 2004; Muchnick and others, 1997; Aho et al., 2006; Cooper and Torczon, 2011; Kennedy and Allen, 2001; Aldrich, 2019b, a).
+Prophecy variables have recently been applied for program verification in a Hoare program logic based on separation logic (Jung et al., 2020).
+They have also been used as a mechanism to support reasoning about transformations between different program representations (Lal and Reps, ; Cook and Koskinen, ).
 Our purpose is different, specifically to use prophecy variables to enable forward reasoning and eliminate the need for backwards analyses and explicit intermediate representations in the context of a lightweight DSL implementation system that uses information about the future execution of the program to optimize the generated code.
@@
 Many DSLs provide domain abstractions directly in the programming language to enhance developer productivity.
-Examples include visualization/graphics languages such as Tikz  (tikz), Vega/Vega-Lite (vega; vegalite), gnuplot (gnuplot), matplotlib (matplotlib), and Graphviz (graphviz); build system and build system generation languages such as make (make), cmake (cmake), and Ninja (ninja); and lexing and parsing languages like flex/lex bison/yacc (flex; yacc) and ANTLR (antlr).
+Examples include visualization/graphics languages such as Tikz  (Tantau, 2013), Vega/Vega-Lite (Contributors, 2025c, b), gnuplot (Williams and Kelley, 2011), matplotlib (Hunter, 2007), and Graphviz (Research, 2025); build system and build system generation languages such as make (Feldman, 1979), cmake (Inc., 2025), and Ninja (Martin, 2012); and lexing and parsing languages like flex/lex bison/yacc (Paxson, 2025; Johnson, 1975) and ANTLR (Parr, 2025).
 Because the main focus of these languages is increased developer productivity and not performance, implementations typically perform little to no forwards or backwards analysis to enable optimizations.
 Our research differs in that it uses prophecy variables for transformations and optimizations that require information about the future execution of the program.
-DSLs such as GraphIt (graphit; g2; ugf), TACO (taco), Halide (halide), Diderot (diderot), Taichi (taichi), SQL (mysql; sqlite) and others target domains in which performance is a primary concern.
+DSLs such as GraphIt (Zhang et al., 2018; Brahmakshatriya et al., 2021b, a), TACO (Kjolstad et al., 2018), Halide (Ragan-Kelley et al., 2013), Diderot (Chiw et al., 2012), Taichi (Hu et al., 2019), SQL (Widenius et al., 2002; Hipp, 2020) and others target domains in which performance is a primary concern.
 Implementations of these DSLs therefore often contain a range of both forward and backward program analyses to enable the generation of optimized code.
@@
 The combination of prophecy variables and reexecution to deliver correct prophecy variable values enables BuildIt to implement transformations and optimizations that require information about the future execution of the program using only forward execution, with no backwards analysis and no intermediate representation.
-Partial evalution/multistage programming is an established language implementation technique with a long history of research (Futamura1971; JonesGomardSestoft1993; ConselDanvy1993; DaviesPfenning1996; TahaSheard1997; TahaSheard2000; CaretteKiselyovShan2009; RompfOdersky2010; Danvy1996; Taha2004).
-Examples of general purpose languages with multistage support include Template Haskell (sheard2002template), Lightweight Modular Staging in Scala (lms), Scheme (scheme), MetaML (TahaSheard2000), Meta OCaml, Mint (mint), Terra (terra; terra2) implemented in Lua, C macros, and C++ templates (vandevoorde2002c++).
+Partial evalution/multistage programming is an established language implementation technique with a long history of research (Futamura, 1971; Jones et al., 1993; Consel and Danvy, 1993; Davies and Pfenning, 1996; Taha and Sheard, 1997, 2000; Carette et al., 2009; Rompf and Odersky, 2010b; Danvy, 1996; Taha, 2004).
+Examples of general purpose languages with multistage support include Template Haskell (Sheard and Jones, 2002), Lightweight Modular Staging in Scala (Rompf and Odersky, 2010a), Scheme (Steele and Sussman, 1975), MetaML (Taha and Sheard, 2000), Meta OCaml, Mint (Westbrook et al., 2010), Terra (DeVito et al., 2013a, b) implemented in Lua, C macros, and C++ templates (Vandevoorde and Josuttis, 2002).
 Multistage programming is often used to embed a DSL within a general purpose language.
-Examples include PyTorch (pytorch), Tensorflow (tensorflow), TVM (Chen2018TVM), AnyDSL (anydsl), Delite (delite), LMS (lms), BuilDSL (buildsl), JAX (jax2018github), SHiM (shim), and BREeze (tamara:meng-thesis:2023).
+Examples include PyTorch (Paszke et al., 2019), Tensorflow (TensorFlow Development Team, 2015), TVM (Chen et al., 2018), AnyDSL (Contributors, 2025a), Delite (Chafi et al., 2011), LMS (Rompf and Odersky, 2010a), BuilDSL (Brahmakshatriya and Amarasinghe, 2022), JAX (Bradbury et al., 2018), SHiM (Ray, 2023), and BREeze (Mitrovska, 2023).
 Such language implementations often include sophisticated program analyses, either via standard programming language implementation techniques, or using history variables or their equivalent (the best of our knowledge, none of these systems use prophecy variables or their equivalent).
@@
 By extending the scope of BuildIt to include languages whose efficient implementation requires these kinds of transformations and optimizations, the presented techniques promise to further eliminate barriers to rapid and efficient DSL implementation, enabling more developers to benefit from these effective application implementation platforms.
-Supplemental Material for Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System
+Supplemental Material for Backwards Data-Flow Analysis using Prophecy Variables in the BuildIt System
 6.1. Einsum DSL Complete Implementation